### v1

# SofT-GRPO: Surpassing Discrete-Token LLM Reinforcement Learning via Gumbel-Reparameterized Soft-Thinking Policy Optimization

Abstract
The soft-thinking paradigm for Large Language Model (LLM) reasoning can outperform the conventional discrete-token Chain-of-Thought (CoT) reasoning in some scenarios, underscoring its research and application value. However, while the discrete-token CoT reasoning pattern can be reinforced through policy optimization algorithms such as group relative policy optimization (GRPO), extending the soft-thinking pattern with Reinforcement Learning (RL) remains challenging. This difficulty stems from the complexities of injecting stochasticity into soft-thinking tokens and updating soft-thinking policies accordingly. As a result, previous attempts to combine soft-thinking with GRPO typically underperform their discrete-token GRPO counterparts. To fully unlock the potential of soft-thinking, this paper presents a novel policy optimization algorithm, SofT-GRPO, to reinforce LLMs under the soft-thinking reasoning pattern. SofT-GRPO injects the Gumbel noise into logits, employs the Gumbel-Softmax technique to avoid soft-thinking tokens outside the pre-trained embedding space, and leverages the reparameterization trick in policy gradient. We conduct experiments across base LLMs ranging from 1.5B to 7B parameters, and results demonstrate that SofT-GRPO enables soft-thinking LLMs to slightly outperform discrete-token GRPO on Pass@1 (+0.13% on average accuracy), while exhibiting a substantial uplift on Pass@32 (+2.19% on average accuracy).  [^2] Codes and weights are available on https://github.com/zz1358m/SofT-GRPO-master
Machine Learning, ICML
(a) Reasoning with Discrete-Token CoT
(b) Reasoning with the Soft-Thinking Pattern
(c) GRPO (shao2024deepseekmath) for Discrete-Token CoT
(d) Existing Work of Soft-Thinking + GRPO (butt2025soft)
(e) SofT-GRPO (Ours) for Reinforcing Soft-Thinking
Figure 1: The soft-thinking pattern (b) passes the expectation of embeddings to the next LLM step (zhang2025soft), which can surpass the conventional discrete-token CoT (a) without any fine-tuning. However, employing the GRPO algorithm (c) will boost the performance of discrete-token CoT, but existing attempts (d) of applying RLVR to soft-thinking derive inferior performances. The proposed SofT-GRPO (e) provides the first valid RLVR algorithm, which can outperform the discrete-token CoT with GRPO.
1 Introduction
Reasoning with large language models (LLMs) has demonstrated impressive versatility across diverse domains (sprague2024cot). However, most existing reasoning methods rely on generating discrete tokens, which may limit their ability to represent certain abstract concepts (zhang2025soft). In pursuit of better expressing abstract ideas and even implementing a possible latent search tree (wu2025llms), zhang2025soft presents the soft-thinking reasoning pattern. The soft-thinking reasoning pattern replaces each discrete token in the chain-of-thought (CoT) with a continuous representation: a weighted sum of d-dimensional token embeddings, computed from their output probabilities, which is then input to the next LLM decoding step. With suitable sampling techniques—such as the Gumbel-Softmax technique (wu2025llms) or the Dirichlet resampling technique (zhuang2025text)—soft-thinking can outperform conventional discrete-token CoT on a wide range of tasks, without requiring any fine-tuning (wu2025llms).
Recently, a growing body of research has focused on Reinforcement Learning with Verifiable Rewards (RLVR) approaches to further enhance the effectiveness of discrete-token CoT reasoning (wang2024reinforcement; liu2025understanding; yu2025dapo). Among these approaches, Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) has emerged as a particularly compelling framework. GRPO operates by sampling groups of CoT trajectories for each query and updating the policy to favor higher-reward trajectories, yielding substantial improvements on various benchmarks. Notably, discrete-token GRPO can consistently outperform the possible advantage brought about by applying the soft-thinking pattern on original LLMs. In contrast, initial attempts to synergize soft-thinking with GRPO typically underperform in comparison to their discrete-token counterparts (butt2025soft). As shown in 1 (b), different from discrete-token reasoning processes, tokens in soft-thinking reasoning are deterministic without randomness. So, this shortfall can be attributed to challenges of 1) introducing controllable stochasticity within soft-thinking tokens for reasoning path exploration, as well as 2): effectively updating the soft-thinking policies by exploiting high-quality samples. These challenges highlight the need for specialized policy optimization algorithms that fully unlock the latent potential of soft-thinking reasoning in LLMs.
To address these challenges and unlock the promise of soft-thinking, we propose a novel policy optimization algorithm, SofT-GRPO, specifically designed to reinforce the soft-thinking reasoning ability in LLMs. As shown in Figure 1(e), in the rollout process, SofT-GRPO samples groups of soft-thinking reasoning paths by injecting sampled Gumbel noises into the output probabilities and employs the Gumbel-Softmax technique to avoid invalid inputs outside the pre-trained discrete-token embedding space. For the following soft-thinking policy updates, SofT-GRPO leverages the reparameterization trick on the Gumbel distribution. It can achieve accurate gradient estimation and accurately attribute the reward improvement to the output probability of LLMs. We conduct thorough evaluations of SofT-GRPO on three representative LLM architectures—DeepSeek-R1-Distill-Qwen-1.5B, LLaMA-3.2-3B-Instruct, and DeepSeek-R1-Distill-Qwen-7B—across five numerical reasoning benchmarks. Experimental results consistently show that SofT-GRPO-enhanced soft-thinking not only slightly surpasses discrete-token GRPO on Pass@1, but also yields substantial gains on Pass@32, highlighting the practical advantages of robust policy optimization for soft-thinking reasoning.
Our contributions can be summarized as follows:
  • We introduce SofT-GRPO, a novel and powerful policy optimization algorithm designed for reinforcing the soft-thinking reasoning paradigm in LLMs. It integrates the Gumbel-Softmax technique into the group rollout process, actively obtaining diverse but valid soft-thinking reasoning paths.
  • We propose an innovative gradient estimation approach via Gumbel reparameterization, enabling precise attribution of improvements to the LLM’s output probability distributions in policy optimization.
  • We conduct comprehensive experiments across LLMs of 1.5B–7B parameters on five in-domain benchmarks and three out-of-domain benchmarks, demonstrating that SofT-GRPO consistently outperforms the discrete-token GRPO baselines, especially at higher sample rates (Pass@16 or Pass@32).
2 Preliminaries
2.1 Discrete-Token CoT Reasoning
Discrete-token CoT reasoning processes seek to solve a $|\boldsymbol{Q}|$-token question $\boldsymbol{Q}=(q_{1},\ldots,q_{|\boldsymbol{Q}|})$ by generating $|\boldsymbol{R}|$ reasoning CoT tokens $\boldsymbol{R}=(r_{1},\ldots,r_{|\boldsymbol{R}|})$ before outputting the answer prediction $\boldsymbol{A}=(a_{1},\ldots,a_{|\boldsymbol{A}|})$ (guo2025deepseek; sprague2024cot). All these tokens are in the discrete language domain, i.e., $\boldsymbol{Q},\boldsymbol{R},\boldsymbol{A}\subset\mathcal{T}$, where $\mathcal{T}$ represents the set of language tokens. Language reasoning tokens and answer tokens are generated with the next-token prediction (NTP) policy of LLM $\pi_{\theta}$ as follows:
$p(\boldsymbol{R},\boldsymbol{A}|\boldsymbol{Q})=$ $\prod_{t=1}^{|\boldsymbol{R}|}\pi_{\theta}(r_{t}|[\boldsymbol{Q},(r_{1},\ldots,r_{t-1})])$ (1)
$\quad\prod_{t=1}^{|\boldsymbol{A}|}\pi_{\theta}(a_{t}|[\boldsymbol{Q},\boldsymbol{R},(a_{1},\ldots,a_{t-1})]),$
where $\cdot$,\ $\cdot$,\ $\cdot$] denotes concatenation.
$\mathcal{J}_{\text{GRPO}}(\theta)=\frac{1}{G}\mathbb{E}_{\boldsymbol{Q}\sim\mathcal{D},\{\boldsymbol{R}\}_{g=1}^{G},\{\boldsymbol{A}\}_{g=1}^{G}\sim p(\cdot,\cdot|\boldsymbol{Q})}$ (2)
$\Bigg[\sum_{g=1}^{G}\frac{1}{\left|\boldsymbol{R}_{g}\right|+\left|\boldsymbol{A}_{g}\right|}\sum_{t=1}^{\left|\boldsymbol{R}_{g}\right|+\left|\boldsymbol{A}_{g}\right|}\Big(\min\left(p_{g,t}\hat{A}_{g},\text{clip}(p_{g,t},1-\epsilon,1+\epsilon)\hat{A}_{g}\right)-\beta D_{KL}(\pi_{\theta}||\pi_{\theta_{\text{ref}}})\Bigg]$
[…]
Supervised Fine-tuning (SFT) for Discrete-Token CoT LLM Reasoning. As straightforward fine-tuning methods for LLM reasoning, SFT methods (wu2025llm; zheng2025reasoning) first collect high-quality CoT labels $\boldsymbol{R}^{*},\boldsymbol{A}^{*}]$ for each question $\boldsymbol{Q}$. Then LLMs are fine-tuned for correct predictions on each of the $|\boldsymbol{R}|+|\boldsymbol{A}|$ tokens. However, SFT methods for LLM reasoning highly rely on the quality of CoT labels, which are difficult to obtain for complex datasets. Moreover, there are also concerns about the out-of-domain generalization ability of SFT (chu2025sft).
RLVR Fine-tuning for Discrete-Token Reasoning. RLVR fine-tuning methods such as GRPO (liu2024deepseek), Dr. GRPO (liu2025understanding), DAPO (yu2025dapo), and Lite PPO (liu2025part) sample several CoTs $\boldsymbol{R},\boldsymbol{A}]$ and assign a reward to each of them based on the quality of the answers $\boldsymbol{A}$. The standard discrete-token GRPO algorithm (shao2024deepseekmath) samples $G$ CoTs for each question $\boldsymbol{Q}$ and optimizes the reasoning policy towards the CoTs with higher rewards.
The loss function of GRPO is shown in Eq. (2), where $\hat{A}_{g}$ represents the advantage function for the $g$-th CoT, the reward function $f(\boldsymbol{A}_{t})=1$ if and only if the answer $\boldsymbol{A}_{t}$ is correct, $\epsilon$ is the clipping hyperparameter, $\pi_{\theta_{\text{old}}}$ is the policy before the update, and $D_{KL}$ represents the KL-divergence of policy $\pi_{\theta}$ and a reference model policy $\pi_{\theta_{\text{ref}}}$. Employing RLVR methods, the average performance of LLMs in numerical reasoning problems can be improved without the need for labels (yue2025does).
2.2 Soft-Thinking Paradigm
Conventional discrete-token reasoning is constrained to selecting a single token from the token set $\mathcal{T}$ at each of the CoT reasoning steps. This approach may hinder the model’s ability to express certain abstract concepts that cannot be easily represented by a single deterministic token (zhang2025soft). To enhance the model’s capability to represent abstract concepts, zhang2025soft presents another paradigm called soft-thinking. The proposed soft-thinking paradigm replaces discrete-token reasoning steps $\boldsymbol{R}$ with a soft-thinking reasoning path $\boldsymbol{S}=(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{|\boldsymbol{S}|})$. Each token $\boldsymbol{s}_{i}\in\mathcal{R}^{d}$ is a real-valued vector, which is calculated as the weighted sum of token embeddings with their output probabilities. The weighted sum embeddings within the vector space of token embeddings are then fed into the next LLM step as follows:
$\boldsymbol{p}_{t}\sim\pi_{\theta}(\cdot|\boldsymbol{Q},(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{t-1})])),$ (3)
$\boldsymbol{s}_{t}=\sum_{i=1}^{|\mathcal{T}|}p_{i}\cdot\boldsymbol{e}_{i},$
where $p_{i}\in[0,1]$ is the predicted probability of token $i$ in $\boldsymbol{p}_{t}$ and $\boldsymbol{e}_{i}\in\mathcal{R}^{d}$ is the LLM embeddings of token $i$. The soft-thinking paradigm may inherently implement a multi-thread latent search tree in the serial decoding process (wu2025llms).
Based on Eq. 3, wu2025llms find that most pre-trained LLMs tend to be single-thread, and wu2025llms propose to introduce sampling strategies in the soft-thinking reasoning process for randomness, employing methods such as the Gumbel-Softmax technique (jang2016categorical) based on the output probability $\boldsymbol{p}_{t}$ as follows:
$g_{i}=\log p_{i}+\epsilon_{i},$ $y_{i}=\dfrac{\exp(g_{i}/\tau_{g})}{\sum_{i=1}^{|\mathcal{T}|}\exp(g_{i}/\tau_{g})},$ (4)
$\boldsymbol{s}_{t}=\sum_{i=1}^{|\mathcal{T}|}y_{i}\cdot\boldsymbol{e}_{i},$
where $\epsilon_{i}$ is a scaler noise sampled from the Gumbel distribution Gumbel(0,1)(0,1), and $\tau_{g}$ is the temperature of Gumbel-Softmax. Besides using Gumbel-Softmax. wu2025llms also tries to use the Dirichlet resampling technique as follows:
$(x_{1},\ldots,x_{|\mathcal{T}|})\sim f(\boldsymbol{p}_{t})=\frac{1}{B(\alpha\cdot\boldsymbol{p}_{t})}\prod_{i}^{|\mathcal{T}|}x_{i}^{\alpha\cdot p_{i}-1}$ (5)
$\qquad\boldsymbol{s}_{t}=\sum_{i=1}^{|\mathcal{T}|}x_{i}\cdot\boldsymbol{e}_{i},$
where $\alpha$ is a scaling parameter. Empirically, cooperated with the Gumbel-Softmax technique for randomness, the soft-thinking pattern can outperform conventional discrete-token CoT on a broad range of tasks, including numerical reasoning, code reasoning, and scientific reasoning, without requiring any fine-tuning. (wu2025llms).
Figure 2: The pipeline of the proposed SofT-GRPO algorithm. In training with a Query $\boldsymbol{Q}$, the SofT-GRPO first generates a group of $G$ soft-thinking reasoning paths with Gumbel noises and the Gumbel-Softmax technique (jang2016categorical). We transmit the value $g^{\prime}_{i}$ and $y^{\prime}_{i}$ for the loss calculation afterward. Then, we reconstruct the soft-thinking input. Finally, we update the soft-thinking policy with the off-policy REINFORCE (williams1992simple) algorithm, optimizing the soft-thinking reasoning tokens with Gumbel reparameterization.
2.3 Attempts of RLVR on Soft-Thinking
However, after being boosted with the RLVR fine-tuning, discrete-token CoT will turn to clearly outperform the soft-thinking reasoning. So, butt2025soft tries to similarly improve the performance of soft-thinking using GRPO. butt2025soft adds Gaussian noise on the input $s_{t}$ in Eq. (3) as follows:
$\hat{\boldsymbol{s}}_{t}=\boldsymbol{s}_{t}+\mathcal{N}(0,\sigma^{2}Id).$ (6)
In deriving the off-policy soft-thinking probability in GRPO, they restore the value of $\hat{\boldsymbol{s}}_{t}$ in the rollout and calculate the log probability for the soft-thinking reasoning path $\boldsymbol{S}$ with the Gaussian reparameterization trick as follows, keeping the updating process of the answering part $\mathbf{A}$ unchanged:
[…]
However, we observe that there are two drawbacks to this method. 1): Although it may help avoid inputs outside the pre-trained discrete-token embedding space, adding noise to inputs instead of logits is not direct and may theoretically mismatch the LLM predictions. As analyzed in Appendix C.1, the embeddings of each token may be linearly dependent, so it becomes hard to attribute which probability $p_{i}$ may contribute to effectiveness. Moreover, the added noise may even be impossible to represent by the embeddings. 2): The methods in (butt2025soft) do not use any advanced sampling methods for effectiveness in training (e.g., incorporating Dirichlet resampling or Gumbel Softmax technique (wu2025llms)), which will undermine the performance of sampled soft-thinking reasoning paths.
So empirically, although observing improvements compared to discrete-token GRPO fine-tuning on the Pass@32 metrics (i.e., the pass rate with 32 attempts), there is a severe degradation on the average accuracy (butt2025soft). As a result, existing attempts of applying GRPO to soft-thinking will not keep its advantage over the discrete-token CoT under the no-finetune setting.
$\mathcal{J}_{\text{SofT-GRPO}}(\theta)=\frac{1}{G}\mathbb{E}_{\boldsymbol{Q}\sim\mathcal{D},\{\boldsymbol{S}\}_{g=1}^{G},\{\boldsymbol{A}\}_{g=1}^{G}\sim p(\cdot,\cdot|\boldsymbol{Q})}$ (8)
$\Bigg[\sum_{g=1}^{G}\frac{1}{\left|\boldsymbol{S}_{t}\right|+\left|\boldsymbol{A}_{t}\right|}\sum_{t=1}^{\left|\boldsymbol{S}_{t}\right|+\left|\boldsymbol{A}_{t}\right|}\Big(\min\left(p_{g,t}\hat{A}_{g},\text{clip}(p_{g,t},1-\epsilon,1+\epsilon)\hat{A}_{g}\right)-\nabla\beta D_{KL}(\pi_{\theta}||\pi_{\theta_{\text{ref}}})\Bigg]$
[…]
3 SofT-GRPO: Reinforcing Soft-Thinking Policy with Gumbel Reparameterization
To fully unlock the potential of the soft-thinking paradigm and establish a reliable framework to reinforce it, we present an effective RLVR algorithm for soft-thinking policies, SofT-GRPO. As shown in Figure 2, the proposed SofT-GRPO first samples soft-thinking CoTs with controllable randomness using the Gumbel-Softmax technique (wu2025llms; jang2016categorical). In the following policy update stage, we propose a novel SofT-GRPO loss function using the Gumbel reparameterization trick.
3.1 Group Rollout with Gumbel Noise
As an off-policy RLVR algorithm, for each Query $\boldsymbol{Q}$, SofT-GRPO samples and restores a group of $G$ soft-thinking CoTs in the rollout stage in parallel. As discussed in Section 2.2, advanced sampling methods will bring performance improvement together with stochasticity. So, to explore diverse and powerful soft-thinking reasoning paths, we introduce the Gumbel-Softmax resampling technique (shown in Eq. (4)) in the rollout process as follows:
$(p_{1},\ldots,p_{|\mathcal{T}|})=\pi_{\theta_{\text{old}}}(\cdot|[\boldsymbol{Q},(s_{1},\ldots,s_{t-1})]),$ (9)
$g^{\prime}_{i}=\log\ p_{i}+\epsilon_{i},\quad y^{\prime}_{i}=\dfrac{\exp(g^{\prime}_{i}/\tau_{g})}{\sum_{i=1}^{|\mathcal{T}|}\exp(g^{\prime}_{i}/\tau_{g})},$
$\boldsymbol{s}_{t}=\sum_{i=1}^{|\mathcal{T}|}y^{\prime}_{i}\cdot\boldsymbol{e}_{i},$
where $\boldsymbol{e}_{i}$ is the d-dimensional embedding of the $i$-th token, $\pi_{\theta_{\text{old}}}$ is the old LLM policy used in rollout, and $\epsilon_{i}$ is the Gumbel noise. Using the inverse transform sampling, we can sample $\epsilon_{i}$ by computing $g=-\log(-\log(u))$ where $u\sim\text{Uniform}(0,1)$ (jang2016categorical). Different from adding noise in other distributions, the Gumbel Softmax technique can ensure the stability of the added Gumbel Noise. As shown in Theorem 10 (Refer to Appendix C.2 for proof), the Gumbel Softmax will preserve the distribution of LLM’s output probabilities after adding noise for stochasticity. So, it may simulate the multinomial sampling and reduce the risk of making $\boldsymbol{s}_{t}$ drop outside the pre-trained discrete-token embedding space.
Theorem 3.1 (Gumbel-max Trick).
Let $(p_{1},\dots,p_{n})$ be nonnegative, and $\epsilon_{1},\dots,\epsilon_{n}$ independent samples from $\mathrm{Gumbel}(0,1)$ (maddison2016concrete),
$\Pr\left(j=\arg\max_{i}\left(\epsilon_{i}+\log p_{i}\right)\right)=\frac{p_{j}}{\sum_{i=1}^{n}p_{i}}.$ (10)
3.2 Gumbel Reparameterization for Loss Function
When fine-tuning with the conventional discrete-token GRPO, as represented in Eq. (1), the probability of sampled trajectories can be easily obtained from the output Categorical distributions. However, calculating such a probability is more difficult in soft-thinking (jain2025learning). In the rollout process, we estimate the probability with the reparameterization trick over Gumbel noises as follows:
$\log p(\boldsymbol{s}_{t}|[\boldsymbol{Q},(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{t-1})],\theta_{\text{old}})=\sum_{i=1}^{|\mathcal{T}|}-\epsilon_{i}-\exp(-\epsilon_{i}),$ (11)
where $\boldsymbol{s}_{t}$ is built from Eq. (9). When doing the on-policy sampling in SofT-GRPO, we reconstruct the group of soft-thinking reasoning paths $\boldsymbol{S}=(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{|\boldsymbol{S}|})$ from the restored $y^{\prime}_{i}$ values in the rollout process in Eq. (9). Then, the log prob is generated with the restored $g^{\prime}_{i}$ values as follows:
$\log p(\boldsymbol{s}_{t}|[\boldsymbol{Q},(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{t-1})],\theta)=$ (12)
$\qquad\quad\sum_{i=1}^{|\mathcal{T}|}-(g^{\prime}_{i}-\log p_{i})-\exp(-(g^{\prime}_{i}-\log p_{i})),$
$\text{where}\quad(p_{1},\ldots,p_{|\mathcal{T}|})=\pi_{\theta}(\cdot|[\boldsymbol{Q},(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{t-1})]).$
Finally, we can represent the total loss of the proposed SofT-GRPO as shown in Eq. (8)
3.3 Detailed Settings
We follow the general setting in soft-thinking works (wu2025llms), enabling both the top-p and top-k sampling strategies. In both training and inference, we set top-p as $0.95$ and top-k as $5$, the temperature of LLMs $\tau=0.6$, and the temperature in Gumbel Softmax $\tau_{g}=0.1$.
Table 1: Experiment results of baselines and the proposed Latent-RPC on five numerical reasoning benchmarks. We cover 3 base LLMs from 1.5B to 7B and two reasoning patterns, i.e., discrete-token CoT and soft-thinking reasoning. @1 metrics denote the Mean@32 values, where we run each method 32 times on the dataset for average Pass@1 accuracies. @16 and @32 denote the Pass@16 and Pass@32 values on the dataset, respectively. We multiply all the results by 100 to highlight the differences between results. The best result on each metric and dataset is underlined, the best average result is bolded, and the second-best average result is shaded.
Dataset AIME2024 AIME2025 AMC23 MATH-500 GSM8K Average
Metrics @1 @16 @32 @1 @16 @32 @1 @16 @32 @1 @16 @32 @1 @16 @32 @1 @16 @32
DeepSeek-R1-Distill-Qwen-1.5B Base LLM
Discrete-Token CoT Reasoning Pattern
No-Finetune 30.6 70.0 73.3 23.0 46.7 53.3 70.7 92.5 95.0 84.6 97.8 97.8 81.5 95.8 96.7 58.09 80.54 83.23
+ GRPO 31.8 66.7 76.7 25.3 46.7 46.7 77.3 95.0 95.0 87.1 97.4 97.8 84.9 95.1 95.8 61.28 80.16 82.39
Soft-Thinking Reasoning Pattern
No-Finetune 27.3 66.7 70.0 23.8 46.7 53.3 69.9 95.0 95.0 79.4 93.2 96.6 81.0 94.6 97.1 56.28 79.23 82.41
+ GRPO 29.2 70.0 73.3 25.4 46.7 53.3 75.8 95.0 95.0 86.3 96.8 98.2 84.9 95.6 96.4 60.31 80.81 83.26
+ SofT-GRPO 32.6 76.7 80.0 26.1 50.0 53.3 76.4 97.5 97.5 86.3 97.4 98.0 85.5 96.1 97.0 61.39 83.54 85.18
LLaMA-3.2-3B-Instruct Base LLM
Discrete-Token CoT Reasoning Pattern
No-Finetune 4.4 20.0 26.7 0.3 0.3 1.0 18.3 65.0 75.0 38.1 75.6 84.0 67.9 92.1 94.6 25.79 50.61 56.26
+ GRPO 7.3 23.3 26.7 0.5 3.3 3.3 27.3 62.5 67.5 48.3 77.2 82.6 79.6 95.4 96.5 32.60 52.35 55.32
Soft-Thinking Reasoning Pattern
No-Finetune 3.4 16.7 16.7 0.2 6.7 6.7 17.6 70.0 77.5 36.7 76.0 81.4 66.9 91.6 94.7 24.96 52.18 55.39
+ GRPO 8.0 20.0 23.3 0.7 3.3 10.0 27.3 70.0 75.0 47.8 76.8 81.8 79.2 94.8 96.3 32.60 53.00 57.28
+ SofT-GRPO 7.7 23.3 26.7 0.3 10.0 10.0 31.3 67.5 67.5 47.2 77.6 83.4 77.6 96.4 97.7 32.83 54.96 57.06
DeepSeek-R1-Distill-Qwen-7B Base LLM
Discrete-Token CoT Reasoning Pattern
No-Finetune 55.7 80.0 80.0 39.4 66.7 66.7 89.9 97.5 97.5 93.6 98.8 99.2 89.5 96.7 97.0 73.62 87.94 88.07
+ GRPO 54.0 80.0 80.0 40.4 66.7 70.0 89.2 95.0 95.0 93.5 98.4 99.0 91.4 96.1 96.6 73.69 87.24 88.12
Soft-Thinking Reasoning Pattern
No-Finetune 55.3 80.0 83.3 39.2 66.7 70.0 90.2 95.0 97.5 93.3 98.8 99.0 89.3 96.6 97.0 73.44 87.41 89.38
+ GRPO 55.5 80.0 80.0 37.8 63.3 66.7 90.1 97.5 97.5 93.6 98.6 99.0 91.8 96.7 96.9 73.76 87.23 88.01
+ SofT-GRPO 53.2 80.0 83.3 40.4 60.0 73.3 89.6 97.5 97.5 93.3 98.6 99.0 92.1 97.2 97.7 73.74 86.66 90.18
4 Experiments
In this section, we implement the proposed SofT-GRPO algorithm to reinforce the soft-thinking reasoning of three LLMs, including DeepSeek-R1-Distill-Qwen-1.5B, LLaMA-3.2-3B-Instruct, and DeepSeek-R1-Distill-Qwen-7B.
4.1 Implementation Detail
Training & Testing Settings
We employ DeepScaler (deepscaler2025) as the training dataset, which contains 40,315 queries. In implementing SofT-GRPO, we use SGLang (zheng2024sglang) [^1] Specifically, we modify the rollout process based on the SGLang implementation in https://github.com/eric-ai-lab/Soft-Thinking for the rollout process and the verl-0.4.x framework (sheng2024hybridflow) for policy update.
We involve five famous numerical benchmarks as test sets, i.e., AIME2024, AIME2025, AMC23, MATH-500, and GSM8K. The maximum generation length is confined to 8192 in training & validating, and 32768 in testing. Test answers are verified using the Math Verify package (Kydlicek_Math-Verify_Math_Verification). All experiments in this paper are implemented on a node of 8× NVIDIA H200 GPUs (141 GB VRAM each). With the 64 batch size and a 1e-6 learning rate, running SofT-GRPO on a 1.5B LLM will take approximately 45 hours.
Table 2: Comparison of the proposed SofT-GRPO to the RLVR fine-tuning method proposed in (butt2025soft). Due to the difficulty in reproduction, we compare to their reported results under the same training dataset. GRPO* represents the reported result of GRPO in its paper, and Soft Tokens* represents its reported results fine-tuned under the soft-thinking pattern.
Dataset MATH-500 GSM8K
Metrics @1 @32 @1 @32
LLaMA-3.2-3B-Instruct Base LLM
Discrete-Token CoT Reasoning Pattern
No-Finetune 38.1 84.0 67.9 94.6
+ GRPO 48.3 82.6 79.6 96.5
+ GRPO* 48.3 78.1 79.5 96.6
+ Soft Tokens* 37.4 80.6 71.0 98.0
Soft-Thinking Reasoning Pattern
+ Soft Tokens* 41.3 77.9 75.5 95.2
47.2 83.4 77.6 97.7
+ SofT-GRPO (+5.9) (+5.5) (+2.1) (+2.5)
Table 3: Average accuracies on out-of-domain datasets. We cover GPQA Diamond, HumanEval, and MBPP. @1 metrics denote the Mean@32 values, where we run the methods 32 times on the dataset for average Pass@1 accuracies. @8, @16, and @32 denote the Pass@8, Pass@16, and Pass@32 values on the dataset, respectively.
Dataset GPQA Diamond HumanEval MBPP Average
Metrics @1 @8 @16 @32 @1 @8 @16 @32 @1 @8 @16 @32 @1 @8 @16 @32
DeepSeek-R1-Distill-Qwen-1.5B Base LLM
Discrete-Token CoT Reasoning Pattern
No-Finetune 36.7 84.3 92.4 96.0 68.1 87.2 90.2 93.9 65.5 84.8 89.1 91.1 56.77 85.45 90.59 93.64
+ GRPO 35.4 77.8 88.4 93.4 72.2 90.9 92.7 94.5 68.1 85.2 87.2 90.0 58.56 84.62 89.41 92.65
Soft-Thinking Reasoning Pattern
No-Finetune 36.0 83.8 91.9 97.0 67.2 89.6 91.5 92.7 64.7 84.4 86.0 88.7 55.98 85.97 89.79 92.79
+ GRPO 36.5 81.3 91.4 94.4 71.8 89.0 92.7 95.1 68.1 84.8 87.9 90.3 58.82 85.05 90.68 93.28
+ SofT-GRPO 37.3 82.8 89.9 95.5 71.2 88.4 91.5 94.5 68.8 88.7 90.3 91.4 59.08 86.65 90.54 93.80
Table 4: Experiments on using majority voting to boost the performances on AIME2024, AIME2025, AMC23 and GSM8K. @1 represents the Pass@1 result from Table 1, which is averaged over 32 runs. M@16 and M@32 represent Major@16 and Major@32, respectively.
Dataset AIME2024 AIME2025 AMC23 GSM8K Average
Metrics @1 M@16 M@32 @1 M@16 M@32 @1 M@16 M@32 @1 M@16 M@32 @1 M@16 M@32
DeepSeek-R1-Distill-Qwen-1.5B Base LLM
Discrete-Token CoT Reasoning Pattern
No Fine-tune 30.6 56.7 60.0 23.0 36.7 40.0 70.7 90.0 95.0 81.5 89.5 89.5 51.5 68.2 71.1
+ GRPO 31.8 53.3 50.0 25.3 36.7 30.0 77.3 92.5 92.5 84.9 90.1 90.4 54.8 68.1 65.7
Soft-Thinking Reasoning Pattern
No Fine-tune 27.3 56.7 60.0 23.8 33.3 33.3 69.9 95.0 92.5 81.0 89.3 89.3 50.5 68.6 68.8
+ GRPO 29.2 43.3 46.7 25.4 30.0 30.0 75.8 92.5 92.5 84.9 90.5 90.8 53.8 64.1 65.0
+ SofT-GRPO 32.6 60.0 63.3 26.1 33.3 36.7 76.4 92.5 95.0 85.5 90.6 90.5 55.2 69.1 71.4
Baselines
We include discrete-token GRPO trained on the same dataset and base LLMs without fine-tuning as baselines. In comparing the results of baselines and the proposed SofT-GRPO, we cover two reasoning patterns, i.e., conventional discrete-token CoT reasoning and soft-thinking reasoning. For baselines trained with discrete tokens, we implement their soft-thinking reasoning pattern with the effective Gumbel-Softmax technique (the temperature $\tau_{g}=0.5$) proposed in wu2025llms. Under both patterns, we follow the hyperparameters in (wu2025llms), setting the temperature $\tau$ of LLMs to 0.6 (we discuss this setting in Appendix D.2), top-p to 0.95, and top-k to 30.
Metrics
In evaluating the performances of LLMs, we cover the general metrics of Mean@32, Pass@16, and Pass@32. Mean@32 judges the average Pass@1 accuracy over 32 runs on the datasets. Pass@16 and Pass@32 measure the average probability of covering the accurate answer within 16 and 32 runs, respectively.
4.2 Main Result
The main result of SofT-GRPO is shown in Table 1. SofT-GRPO can lead to a clear and consistent improvement from the No-Finetune results under the soft-thinking reasoning pattern. Compared to GRPO under the discrete-token CoT, the proposed SofT-GRPO exhibits slight but stable leads (+0.13% on average) on the Mean@32 (i.e., Pass@1, @1 in the Table) metrics over LLMs of three sizes. Notably, the SofT-GRPO can lead to a clear improvement in the Pass@16 (+1.80% on average) and Pass@32 (+2.19%) metrics. Moreover, this observation will still hold when adopting the soft-thinking pattern on the GRPO, demonstrating the significance of SofT-GRPO in reinforcing the soft-thinking performance.
4.3 Comparison to butt2025soft
In this subsection, we conduct comparison experiments with an existing RLVR algorithm for the soft-thinking pattern, the method in butt2025soft. We note this method as Soft Tokens in Table 2, where the proposed SofT-GRPO can demonstrate a clear update compared to the results reported in butt2025soft. Compared to SofT-GRPO, the algorithm in butt2025soft requires transitioning the $d$-dimensional vector between rollout workers and RLVR workers, improving the difficulty of implementation.
Table 5: Ablation studies on the noise added in the proposed SofT-GRPO. We highlight the performance differences on the ablation variants (i.e., adding Dirichlet noises or Gaussian noises).
Dataset AIME2024 AIME2025 AMC23 MATH-500 GSM8K Average
Metrics @1 @16 @32 @1 @16 @32 @1 @16 @32 @1 @16 @32 @1 @16 @32 @1 @16 @32
DeepSeek-R1-Distill-Qwen-1.5B Base LLM
Soft-Thinking Reasoning Pattern
No-Finetune 27.3 66.7 70.0 23.8 46.7 53.3 69.9 95.0 95.0 79.4 93.2 96.6 81.0 94.6 97.1 56.3 79.2 82.4
+ GRPO 29.2 70.0 73.3 25.4 46.7 53.3 75.8 95.0 95.0 86.3 96.8 98.2 84.9 95.6 96.4 60.3 80.8 83.3
+ SofT-GRPO (Original, Gumbel) 32.6 76.7 80.0 26.1 50.0 53.3 76.4 97.5 97.5 86.3 97.4 98.0 85.5 96.1 97.0 61.4 83.5 85.2
25.4 60.0 70.0 22.9 56.7 63.3 68.4 97.5 97.5 83.6 97.0 97.8 83.6 97.0 97.8 56.8 81.6 85.3
+ SofT-GRPO (Dirichlet Noise) (-7.2) (-16.7) (-10) (-3.2) (+6.7) (+10) (-8) (0) (0) (-2.7) (-0.4) (-0.2) (-1.8) (+0.9) (+0.8) (-4.6) (-1.9) (+0.1)
21.9 60.0 73.0 20.0 40.0 43.3 65.6 97.5 97.5 79.9 95.2 96.6 81.8 94.5 95.5 53.8 77.4 81.2
+ SofT-GRPO (Gaussian Noise) (-10.7) (-16.7) (-7) (-6.1) (-10) (-10) (-10.8) (0) (0) (-6.5) (-2.2) (-1.4) (-3.7) (-1.7) (-1.5) (-7.6) (-6.1) (-4)
(a) Training reward curve of variants on added noises
(b) Validation reward curve of different hyper-parameters
(c) Training reward curve of different hyper-parameters
Figure 3: Smoothed training or validation curves of ablation studies (the dashed background contains the actual data points). (a) discusses the setting of adding Gumbel noise in SofT-GRPO. (b) discusses the setting of top-p=0.95 and the Gumbel-Softmax temperature $\tau_{g}=0.1$ in SofT-GRPO.
4.4 Comparison in Out-of-Domain Datasets
Besides numerical reasoning, we conduct out-of-domain experiments to evaluate the general reasoning ability of SofT-GRPO. As shown in Table 3, the LLMs fine-tuned with SofT-GRPO on numerical queries can still demonstrate advantages from Pass@1 to Pass@32 on a scientific reasoning benchmark (GPQA Diamond) and two code benchmarks (HumanEval and MBPP).
4.5 Comparison in the Token Efficiency
In addition to the accuracy, the token efficiency of LLMs is also a crucial metric. In Table 6 and Appendix D.1, we compare the token efficiency of baselines and SofT-GRPO. Compared to the No-Finetune results, fine-tuning with SofT-GRPO will effectively reduce the thinking length. Compared to the discrete-token GRPO, SofT-GRPO will not severely increase the number of tokens. Specifically, there is a clear reduction in the thinking length of the LLaMA-3.2-3B-Instruct model, which demonstrates the effectiveness of SofT-GRPO in saving computational consumption.
4.6 Boosting SofT-GRPO with Majority Voting
As demonstrated in Table 1 and Table 3, SofT-GRPO can lead to clear advantages on Pass@32. To further exploit this advantage, in this subsection, we design to boost SofT-GRPO with majority voting (chen2024more). As shown in Table 4, SofT-GRPO with majority-voting can outperform No-Finetune LLM and LLM fine-tuned with discrete-token GRPO across Major@16 (the accuracy of the most common answer in 16 runs) and Major@32. Results exhibit the potential to strengthen the SofT-GRPO-fine-tuned LLMs for better soft-thinking solvers.
5 Discussion
5.1 Ablation on Noise Added
In SofT-GRPO, we employ the Gumbel Softmax technique for controllable stochasticity. As discussed in Section 2.2, adding Dirichlet noise to predicted probabilities can be another choice, so we compare the performance of the original SofT-GRPO and two of its variants, adding Dirichlet noise or Gaussian noise to probabilities in training. The reward curve in training is shown in Figure 3(a), the validation curve is shown in Figure 3(b), and the final performance of variants is shown in Table 5. LLMs can not learn a clear refinement from the added Dirichlet noise, and adding the Gaussian noise will cause a poorer initial performance. So, these variants cannot perform as well as integrating the original Gumbel noises.
5.2 Ablation on Hyper-Parameters
SofT-GRPO sets the top-p as 0.95 and the Gumbel temperature as $\tau_{g}=0.1$. To demonstrate the reason for these settings, in Figure 3(c), we compare the training reward curve of the original SofT-GRPO with two variants (varying top-p to 1.0 or $\tau_{g}$ to 0.25). Results show that both variants will cause a collapse in the training process. As analyzed in Appendix D.3, the two variants will cause a substantial improvement in the KL divergence between $\pi_{\theta}$ and $\pi_{\theta_{ref}}$. So, we attribute this kind of collapse to the fact that fine-tuning under the soft-thinking pattern may lead to inputs outside the pre-trained discrete-token embedding space.
6 Conclusion
This paper presents a powerful RLVR algorithm, SofT-GRPO, to reinforce LLMs under the soft-thinking reasoning pattern. It integrates controllable stochasticity with Gumbel-Softmax and reinforces the soft-thinking policy with Gumbel reparameterization. The proposed SofT-GRPO can demonstrate better numerical, scientific, and code reasoning ability compared to the conventional discrete-token GRPO on Pass@1, especially Pass@16 and Pass@32. It can also be boosted to a superior solver with majority voting. This article emphasizes the prospects of soft-thinking in LLM reasoning. In the future, we will try building effective tricks over SofT-GRPO and apply it to border fields such as Vision Language Models.
Appendix A Related Work
In this section, we will discuss recent developments on discrete-token RLVR and a similar domain of soft-thinking, the latent reasoning methods.
A.1 RLVR on Discrete-Token CoT Reasoning
The Deepseek-R1-zero (guo2025deepseek) has demonstrated remarkable performance with discrete-token RLVR. And there has been a wide collection of RLVR methods for discrete-token CoT policy optimization, e.g., GRPO (shao2024deepseekmath), Dr. GRPO (liu2025understanding), DAPO (yu2025dapo), and Lite PPO (liu2025part). To further improve these algorithms towards some specific goals, recent works focus on developing RLVR methods for efficient responses (feng2025efficient) or controlling the entropy in CoT generation for better exploration-exploitation balance (hao2025rethinking).
To generate concise discrete-token CoTs, arora2025training modifies the reward function by adding penalties on the generation length. qi2025optimizing and dai2025s turn to formulate the fine-tuning process for concise language CoT as a multi-objective optimization task with the trade-off between token efficiency and accuracy. They truncate the generation of discrete-token CoT at several thinking budgets and optimize the overall performance over them.
RLVR methods with entropy control try to make a better balance between exploration and exploitation. They try to control the entropy in the training process to stimulate the exploration within the groups of CoTs in GRPO wang2025arbitrary.
In the future, we plan to have a comprehensive investigation on whether these modifications can also boost the proposed SofT-GRPO.
A.2 Latent Reasoning
Similar to the soft-thinking pattern, latent reasoning methods pass continuous vectors between LLM steps. These methods fully decouple the reasoning process from explicit natural language (which soft-thinking does not do) and perform inference in the hidden space of the model. Generally, latent CoT methods are often diverse from each other and can be mainly divided into token-wise auto-aggressive methods and auxiliary strategies (chen2025latentcot; tan2025think). Token-wise auto-aggressive methods transform the reasoning process into ’soft thoughts’ with dense latent embeddings (hao2024training) or specialized tokens (e.g., pause (goyal2024pause; zelikman2024quietstar)). These methods focus on transferring the original reasoning policy in the language domain to a latent embedding space, including curriculum learning (e.g., Coconut (hao2024training), LightThinker (zhang2025lightthinker), SIM-COT (wei2025sim)), self-distillation (e.g., CODI (shen2025codi)), and one-shot compression (e.g., CoLaR (tan2025think), SynAdapt (wang2025synadapt)). Auxiliary strategies (e.g., SoftCoT (xu2025softcot)) generate latent embeddings from an auxiliary module and inject them into the frozen main model (cheng2024compressed; xu2025softcot++).
Due to the existing token-wise auto-aggressive methods completely treating the language CoTs as the label, these methods can hardly surpass or even reach the performance level of language LRMs. So empirically, these methods can effectively improve the token efficiency compared to language CoT, but there is a clear performance drop. Auxiliary strategies, instead, can effectively prompt the performance of the original LLM, sacrificing the running efficiency.
Compared to latent reasoning methods mainly aiming at better efficiency, the proposed SofT-GRPO aims at reinforcing the accuracy of the soft-thinking pattern to surpass discrete-token CoT with GRPO on general reasoning tasks.
Appendix B Prompt of Language Reasoning and Latent Generation
In this part, we show the prompt we adopt for reasoning problems, including the in-domain numerical reasoning, out-of-domain GPQA reasoning, and code reasoning. We inherit the prompts in zhang2025soft for out-of-domain benchmarks. We use the same prompt for both the soft-thinking reasoning pattern and the discrete-token reasoning pattern.
[…]
Prompt for GPQA Reasoning. user Please solve the following multiple-choice question. Please show your choice in the answer field with only the choice letter, e.g.,”answer”: ”C”. {Question}
Prompt for Code Reasoning (HumanEval). user Please solve the programming task below in Python. Code should be wrapped in a markdown code block. “‘python {Question} “‘
Prompt for Code Reasoning (MBPP). user Please solve the programming task with test cases below in Python. Make sure your code satisfies the following requirements: 1. The function name and signature must match exactly as specified in the test cases. 2. Your code should be wrapped in a markdown code block without including any test cases. Task: {Question} Test Cases: “‘python {TestCases} “‘
The blue part represents the specific question (query $\boldsymbol{Q}$), the brown part represents the possible test cases provided in the MBPP code reasoning benchmark. We use the code provided in zhang2025soft in the verification process of the responses.
Appendix C Motivation and Theoretical Proof
C.1 Motivation: Mismatch in butt2025soft
butt2025soft adds Gaussian noise on the soft-token inputs as follows:
$\boldsymbol{s}_{t}=\sum_{i=1}^{|\mathcal{T}|}p_{i}\cdot\boldsymbol{e}_{i},\quad\hat{\boldsymbol{s}}_{t}=\boldsymbol{s}_{t}+\mathcal{N}(0,\sigma^{2}Id)$ (13)
and calculates the log probability as follows:
[…]
Let $\boldsymbol{E}\in\mathbb{R}^{|\mathcal{T}|\times d}$ be the embedding matrix, and each token probability vector $\boldsymbol{p}\in\Delta^{|\mathcal{T}|-1}$ corresponds to the soft input $\boldsymbol{s}=\boldsymbol{E}^{\top}\boldsymbol{p}\in\mathbb{R}^{d}$.
Assume the observed noisy soft input is $\hat{\boldsymbol{s}}$, and we define a likelihood
$\log p(\hat{\boldsymbol{s}}\mid\boldsymbol{p})\propto-\frac{1}{2\sigma^{2}}\|\hat{\boldsymbol{s}}-\boldsymbol{E}^{\top}\boldsymbol{p}\|_{2}^{2}.$ (15)
Suppose we want to regard this as a likelihood on $\boldsymbol{p}$. In general, the mapping $\boldsymbol{p}\mapsto\boldsymbol{E}^{\top}\boldsymbol{p}$ is many-to-one: since $|\mathcal{T}|>d$, the kernel of $\boldsymbol{E}^{\top}$ is nontrivial, so
$\exists\,\boldsymbol{p}_{1}\neq\boldsymbol{p}_{2},\quad\text{with}\quad\boldsymbol{E}^{\top}\boldsymbol{p}_{1}=\boldsymbol{E}^{\top}\boldsymbol{p}_{2}$ (16)
Thus, the same $\boldsymbol{s}$ may correspond to infinitely many $\boldsymbol{p}$. This means that, under this Gaussian model, two different token mixtures can lead to the exact same log-probability value. The information about the original token distribution is partially lost in the embedding projection, unless $\boldsymbol{E}$ is invertible (which it is not). Thus, the use of a Gaussian noise model on the embedding space gives a mismatch to the true simplex-based probability geometry.
In summary of Drawback 1: Due to the non-injectivity (non-invertibility) of the embedding transformation, the model $\log p(\hat{\boldsymbol{s}}_{t})\propto-\|\hat{\boldsymbol{s}}_{t}-\boldsymbol{s}_{t}\|^{2}$ does not define a true likelihood on the simplex of token probabilities.
The above mismatch is not only due to the non-invertibility of the embedding matrix. Even if we restrict $\boldsymbol{p}$ to be sparse (nonzero only on a top-$k$ set, which is a general setting of LLMs or a common nature of LLM predictions), and even if the corresponding submatrix $\boldsymbol{E}_{\mathcal{K}}$ is invertible, the process of adding Gaussian noise in the $d$-dimensional embedding space fundamentally breaks the connection to sparse token distributions.
More specifically, after adding Gaussian noise, with a general top-k setting (e.g., kk=10 to 30, and $k<<d$) the perturbed embedding $\hat{\boldsymbol{s}}_{t}=\boldsymbol{s}_{t}+\boldsymbol{\epsilon}$ (with $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\sigma^{2}\boldsymbol{I}_{d})$) will almost surely not lie in the convex hull of any set of $k$ token embeddings. In other words,
$\forall~\hat{\boldsymbol{s}}_{t},~\text{for almost all }\boldsymbol{\epsilon},~\nexists~k\text{-sparse }\boldsymbol{p}\text{ such that }\hat{\boldsymbol{s}}_{t}=\boldsymbol{E}^{\top}\boldsymbol{p}.$ (17)
The set of all top-$k$ soft-token embeddings forms a low-dimensional union of simplices in $\mathbb{R}^{d}$, which is a measure-zero subset of the space. The probability of a randomly perturbed embedding $\hat{\boldsymbol{s}}_{t}$ coinciding with a legal top-$k$ mixture is thus zero.
Therefore, defining the likelihood $p(\hat{\boldsymbol{s}}_{t})$ as a simple Gaussian on embedding space cannot be interpreted as a likelihood on the space of top-$k$ soft tokens—not only due to non-invertibility or nonlinearity, but more fundamentally because most $\hat{\boldsymbol{s}}_{t}$ produced by noise are not realizable by any top-$k$ soft-token distribution.
In summary of Drawback 2: Under the general top-k setting, the likelihood $p(\hat{\boldsymbol{s}}_{t})$ as a simple Gaussian on the embedding space cannot be interpreted as a likelihood on the space of top-$k$ embeddings.
C.2 Proof for Theorem 3.1
Theorem 10 (Gumbel-max Trick) Let $(p_{1},\dots,p_{n})$ be nonnegative real numbers, not all zero. Let $g_{1},\dots,g_{n}$ be independent samples from $\mathrm{Gumbel}(0,1)$. Then,
$\Pr\left(j=\arg\max_{i}(g_{i}+\log p_{i})\right)=\frac{p_{j}}{\sum_{i=1}^{n}p_{i}}.$ (18)
Proof.
For any $j\in\{1,\dots,n\}$,
$\Pr\left(j=\arg\max_{i}(g_{i}+\log p_{i})\right)$ $=\Pr\left(g_{j}+\log p_{j}\geq g_{i}+\log p_{i},\,\,\forall i\neq j\right)$
$=\int_{-\infty}^{+\infty}\prod_{i\neq j}\Pr\left(g_{i}\leq g_{j}+\log p_{j}-\log p_{i}\right)f_{g_{j}}(g_{j})\,dg_{j},$
where $f_{g_{j}}(g)=e^{-g-\exp(-g)}$ is the PDF of the standard Gumbel distribution.
$\Pr(g_{i}\leq t)=F_{\mathrm{Gumbel}}(t)=\exp\left(-e^{-t}\right)$, so
$\prod_{i\neq j}\Pr\left(g_{i}\leq g_{j}+\log p_{j}-\log p_{i}\right)$
$=\prod_{i\neq j}\exp\left(-e^{-(g_{j}+\log p_{j}-\log p_{i})}\right)$
$=\exp\left(-\sum_{i\neq j}e^{-(g_{j}+\log p_{j}-\log p_{i})}\right)$
$=\exp\left(-e^{-g_{j}}\sum_{i\neq j}\frac{p_{i}}{p_{j}}\right).$
The total probability is
$\int_{-\infty}^{+\infty}e^{-g_{j}}e^{-e^{-g_{j}}}\exp\left(-e^{-g_{j}}\sum_{i\neq j}\frac{p_{i}}{p_{j}}\right)dg_{j}$
$=\int_{-\infty}^{+\infty}e^{-g_{j}}\exp\left(-e^{-g_{j}}\left[1+\sum_{i\neq j}\frac{p_{i}}{p_{j}}\right]\right)dg_{j}.$
Let $S=1+\sum_{i\neq j}\frac{p_{i}}{p_{j}}=\frac{\sum_{i=1}^{n}p_{i}}{p_{j}}$. Substitute $y=e^{-g_{j}},\,dg_{j}=-\frac{dy}{y},\,y\in(0,+\infty)$,
$=\int_{y=+\infty}^{0}y\exp(-yS)\left(-\frac{dy}{y}\right)=\int_{0}^{+\infty}\exp(-yS)dy=\frac{1}{S}=\frac{p_{j}}{\sum_{i=1}^{n}p_{i}}.$
Thus, combining all the above equations,
$\Pr\left(j=\arg\max_{i}(g_{i}+\log p_{i})\right)=\frac{p_{j}}{\sum_{i=1}^{n}p_{i}}.$
Appendix D Supplementary of Experiments
Table 6: Experiments on the token efficiency for baselines and the proposed SofT-GRPO. #Token values in Table represent the number of tokens across all queries, and the #Token_c values represent the number of tokens across correct queries.
Dataset AIME2024 AIME2025 AMC23 MATH-500 GSM8K Average
Metrics #Token #Token_c #Token #Token_c #Token #Token_c #Token #Token_c #Token #Token_c #Token #Token_c
DeepSeek-R1-Distill-Qwen-1.5B Base LLM
Discrete-Token CoT Reasoning Pattern
No-Finetune 16241.6 14997.8 16416.3 13448.8 10052.2 9394.2 5616.6 5368.1 1839.3 1772.5 10033.2 8996.3
+ GRPO 8927.6 8535.6 8039.4 6414.2 5001.3 4672.8 3106.1 2958.5 1417.1 1370.6 5298.3 4790.3
Soft-Thinking Reasoning Pattern
No-Finetune 17857.8 16191.3 17569.4 14582.4 11269.9 10482.0 4015.9 3888.3 1699.3 1649.9 10482.5 9358.8
+ GRPO 9383.2 8934.5 8007.2 7325.8 5203.7 4894.4 3233.6 3131.7 1385.5 1349.9 5442.7 5127.2
+ SofT-GRPO 11039.6 10756.1 10519.6 7831.3 5900.2 5630.4 3549.5 3399.2 1577.6 1542.5 6517.3 5831.9
LLaMA-3.2-3B-Instruct Base LLM
Discrete-Token CoT Reasoning Pattern
No-Finetune 5010.5 3334.1 4297.6 2430.9 2790.7 2398.1 1850.1 1494.2 236.6 224.6 2837.1 1976.4
+ GRPO 8991.8 6368.9 8443.8 11088.3 4998.1 3978.5 3334.8 2463.0 502.0 453.6 5254.1 4870.5
Soft-Thinking Reasoning Pattern
No-Finetune 4859.2 2727.7 4934.8 3735.2 3259.9 2624.0 2086.1 1494.1 243.5 227.8 3076.7 2161.8
+ GRPO 9749.0 7627.8 8498.9 10041.4 5389.9 4987.3 3701.9 2764.6 541.9 505.8 5576.3 5185.4
+ SofT-GRPO 829.7 862.9 893.1 879.6 911.5 927.9 632.8 575.7 294.6 292.9 712.3 707.8
DeepSeek-R1-Distill-Qwen-7B Base LLM
Discrete-Token CoT Reasoning Pattern
No-Finetune 13120.7 11511.4 14347.7 11750.7 6346.0 5987.1 3998.9 3939.0 1061.1 1032.0 7774.9 6844.0
+ GRPO 7795.9 7116.1 8003.5 7369.8 4050.9 3719.7 2473.3 2405.5 1146.7 1118.7 4694.0 4346.0
Soft-Thinking Reasoning Pattern
No-Finetune 13017.4 11888.6 14116.8 11507.4 6346.7 6043.6 3947.4 3858.6 996.3 962.2 7684.9 6852.1
+ GRPO 7464.9 6531.4 8291.6 7160.4 3931.7 3736.7 2473.2 2423.0 1117.0 1104.5 4655.7 4191.2
+ SofT-GRPO 8035.8 7556.8 8381.7 7873.7 4008.0 3843.5 2630.2 2583.5 1293.2 1276.4 4869.8 4626.8
D.1 Supplementary of Token Efficiency
As shown in Section 4.5, besides the performance, the token efficiency of LLMs is also an important metric. In this section, we compare the token efficiency of baselines and SofT-GRPO in Table 6. Compared to No-Finetune variants, SofT-GRPO can demonstrate a clear refinement in both the token efficiency across all queries and the token efficiency across correct queries. Compared to discrete-token GRPO, SofT-GRPO will not cause severe token improvement. Specifically, we observe a severe reduction in the thinking length of the LLaMA-3.2-3B-Instruct model. As shown in Figure 4, unlike GRPO, SofT-GRPO maintains and even enhances the token efficiency compared to the base LLM when training progresses, demonstrating the effectiveness of SofT-GRPO in reducing computational consumption.
Figure 4: Token consumption curve on LLaMA-3.2-3B-Instruct Base LLM during training.
D.2 Experiments under Different Temperatures
In Section 4, we adopt the setting of $\tau=0.6$ for discrete-token CoT. To investigate whether other temperature settings will break our observations in our main experiments, similar to yue2025does, we try temperatures from $\tau=0.6$, $\tau=0.8$, $\tau=1.0$, $\tau=1.2$, and $\tau=1.4$.
As shown in Figure 5, we conduct experiments on the 1.5B LLMs (i.e., DeepSeek-R1-Distill-Qwen-1.5B Base LLM) for their average Pass@k accuracies across five numerical reasoning benchmarks (i.e., AIME2024, AIME2025, AMC23, MATH-500, GSM8K), where the proposed SofT-GRPO can demonstrate outstanding results compared to GRPO and No-Finetune variants with various temperatures, from Pass@1 to Pass@32.
Figure 5: Running discrete-token CoT methods (GRPO and No-finetune) with more temperature options on DeepSeek-R1-Distill-Qwen-1.5B Base LLM. Pass@k represents the pass rate within at most k runs, and Pass@1 is additionally averaged from 32 runs. Experiments are run on the five datasets in Table 1 for the average.
D.3 Supplementary of Ablation on Hyper-Parameters
(a) Training KL Divergence curve between $\pi_{\theta_{\text{ref}}}$ and $\pi_{\theta}$
(b) Training Proximal Policy Optimization (PPO) (schulman2017proximal) KL Divergence curve between $\pi_{\theta_{\text{old}}}$ and $\pi_{\theta}$
Figure 6: KL Divergence curves in SofT-GRPO.
In this subsection, we have a further investigation into the hyperparameter settings. As briefly shown in Section 5.2, adopting a higher top-p or $\tau_{g}$ will cause collapses in training. We attribute these collapses to the case that some soft-thinking tokens may become incomprehensible for LLMs.
As shown in Figure 6(a), the variants (varying top-p to 1.0 or varying $\tau_{g}$ to 0.25) will demonstrate higher divergence between the fixed pre-trained $\pi_{\theta_{\text{ref}}}$ and the current policy, which can be an indicator of the inputs outside the pre-trained embedding space.
Recently, qi2025defeating provides excellent insight into the collapse situation in the RLVR fine-tuning. When collapse is caused by a precision issue, they observe a super high KL divergence between $\pi_{\theta_{\text{old}}}$ and $\pi_{\theta}$ (Refer to Figure 3 in (qi2025defeating), $10^{-3}$ even higher). However, in the variants shown in Figure 6(b), we find that their KL divergence in PPO policies is less than $10^{-5}$, indicating that the variants of SofT-GRPO are less likely to meet precision issues.
Appendix E Baselines & Datasets & Licenses
In our experiments, we evaluate and compare the following model baselines and datasets. For each, we detail the official website and usage license.
E.1 Baselines
We mainly include No-Finetune base LLMs, Discrete-Token GRPO, and the method in butt2025soft (noted Soft Token) as baselines.
Base LLMs
This paper includes DeepSeek-R1-Distill-Qwen-1.5B, LLaMA-3.2-3B-Instruct, and DeepSeek-R1-Distill-Qwen-7B as base LLMs.
Discrete-Token GRPO
We utilize the latest verl (sheng2024hybridflow) https://github.com/volcengine/verl/tree/main and vLLM rollout to implement discrete-token GRPO with default parameters.
butt2025soft Soft Token
Due to the requirement of passing d-dimensional inputs $\hat{\boldsymbol{s}}$ between the rollout workers and the verl policy optimization workers, implementing this algorithm requires a high amount of communication between the rollout workers and the policy update workers. So, we report the results in (butt2025soft) for comparison in Table 2.
E.2 Datasets
This paper covers five in-domain numerical reasoning datasets (i.e., AIME2024, AIME2025, AMC23, MATH-500, and GSM8K (cobbe2021training)), one out-of-domain scientific reasoning dataset GPQA-Diamond (rein2024gpqa), and two out-of-domain code reasoning datasets (i.e., HumanEval and MBPP). These datasets are provided in https://github.com/eric-ai-lab/Soft-Thinking. Specifically, due to their answer can be correct with various equivalent forms, we exclude the MATH-500 benchmark for Table 4.
E.3 Licenses
For Base LLM, Dataset, and frameworks, we list their Licenses in Table 7.
Table 7: A summary of licenses.
Resources Type License URL
DeepSeek-R1-Distill-Qwen-1.5B Base LLM MIT License https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Qwen-1.5B
LLaMA-3.2-3B-Instruct Base LLM Llama 3.2 License https://huggingface.co/meta-llama/Llama-3.2-3B-Instruct
DeepSeek-R1-Distill-Qwen-7B Base LLM MIT License https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Qwen-7B
verl RL-framework Apache-2.0 license https://github.com/volcengine/verl
verl-0.4.x RL-framework Apache-2.0 license https://github.com/volcengine/verl/tree/v0.4.x
SGLang for soft-thinking Inference-framework MIT License https://github.com/eric-ai-lab/Soft-Thinking
AIME2024, AIME2025, GSM8K Dataset MIT License https://github.com/eric-ai-lab/Soft-Thinking
AMC23, MATH-500 Dataset Available Online https://github.com/eric-ai-lab/Soft-Thinking
HumanEval, GPQA-Diamond Dataset MIT License https://github.com/eric-ai-lab/Soft-Thinking
MBPP Dataset Apache-2.0 license https://github.com/eric-ai-lab/Soft-Thinking