### v2

# SofT-GRPO: Surpassing Discrete-Token LLM Reinforcement Learning via Gumbel-Reparameterized Soft-Thinking Policy Optimization

Abstract
The soft-thinking reasoning paradigm has demonstrated superior performance over traditional discrete-token Chain-of-Thought (CoT) reasoning in various scenarios. However, while discrete-token CoT reasoning can be reinforced through advanced reinforcement learning with verifiable rewards (RLVR) techniques such as group relative policy optimization (GRPO), extending the soft-thinking reasoning with such strong techniques remains challenging. This difficulty stems from the complexities of injecting stochasticity into soft-thinking tokens and updating soft-thinking policies accordingly. As a result, previous attempts to combine soft-thinking with RLVR typically underperform their discrete-token RLVR counterparts. To fully unlock the potential of soft-thinking, this paper presents a powerful policy optimization algorithm, SofT-GRPO. It injects the Gumbel noise into token probabilities with Gumbel-Softmax for controllable stochasticity, and leverages the Gumbel reparameterization trick to achieve accurate credit assignment to LLM soft-thinking policies. We conduct experiments over LLMs ranging from 1.5B to 7B parameters, where SofT-GRPO enables LLMs with soft-thinking to slightly outperform discrete-token CoT GRPO on Pass@1 (+0.13% on average accuracy), and brings a substantial uplift on Pass@32 (+2.19% on average) [^2] The codes are available at https://github.com/zz1358m/SofT-GRPO-master.
Machine Learning, ICML
(a) LLM Reasoning with Discrete-Token CoT
(b) LLM Reasoning with Soft-Thinking (Zhang et al., 2025b)
(c) GRPO (Shao et al., 2024) for Discrete-Token CoT
(d) Existing Work of Soft-Thinking + GRPO (Butt et al., 2025)
(e) SofT-GRPO (Ours) for Reinforcing Soft-Thinking
Figure 1: The soft-thinking pattern (b) passes the expectation of token embeddings to the next LLM step (Zhang et al., 2025b), which can surpass the conventional discrete-token CoT (a) without any fine-tuning. However, employing the GRPO algorithm (c) will boost the performance of discrete-token CoT over soft-thinking, and existing attempts (d) of applying RLVR to soft-thinking derive inferior performances. The proposed SofT-GRPO (e) provides the first powerful RLVR algorithm, which can outperform the discrete-token CoT with GRPO on Pass@1, especially Pass@K.
1 Introduction
Reasoning with large language models (LLMs) has demonstrated impressive versatility in diverse domains (Sprague et al., 2024). However, most existing reasoning methods rely on the generation of discrete language tokens, which will limit their ability to represent certain abstract concepts that cannot be accurately represented by a discrete token (Zhang et al., 2025b). In pursuit of better expression of abstract ideas, Zhang et al. (2025b) presents the soft-thinking reasoning pattern. As shown in Figure 1 (b), it replaces each discrete token in the chain-of-thought (CoT) with a continuous representation: the weighted sum of $d$-dimensional token embeddings, taking the output probabilities as the weights. With suitable sampling techniques—such as with the Gumbel-Softmax technique (Wu et al., 2025a) or the Dirichlet resampling technique (Zhuang et al., 2025)—soft-thinking can outperform discrete-token CoT on a wide range of tasks without requiring any fine-tuning (Wu et al., 2025a).
Recently, reinforcement learning with verifiable rewards (RLVR) approaches have become standard for improving discrete-token CoT reasoning (Wang et al., 2024; Liu et al., 2025b; Yu et al., 2025). The most representative RLVR method, Group Relative Policy Optimization (GRPO; Figure 1 (c)) (Shao et al., 2024), samples groups of CoT trajectories per query and reinforces the policy toward higher-reward examples. Notably, RLVR fine-tuning can yield discrete-token CoT performance that surpasses that of soft-thinking reasoning from the original, unfine-tuned LLMs. Moreover, after discrete-token GRPO, the model’s soft-thinking ability often does not improve as substantially as its discrete-token reasoning ability.
However, extending RLVR methods to soft-thinking reasoning remains challenging (Jain and Rappazzo, 2025). As illustrated in Figure 1 (b), in contrast to the stochastic nature of sampling-based discrete-token CoTs, soft-thinking reasoning produces deterministic soft-tokens, lacking both randomness and gradients for RLVR. So, extending RLVR to soft-thinking should (1) introduce controllable stochasticity to adequately explore diverse soft-thinking paths, and (2) enable unbiased policy updates that properly assign the credits of high-reward soft-thinking samples to LLM logits. Previous work (Butt et al., 2025) (see Figure 1 (d)) attempts to address stochasticity by injecting Gaussian noise into soft tokens and applying a Gaussian reparameterization trick for RLVR gradients, but such approaches have struggled to match the performance of discrete-token RLVR, largely due to difficulties in attributing RLVR advantages to logits.
To address these challenges and fully unlock the LLM reasoning ability by soft-thinking, we propose a powerful policy optimization algorithm, SofT-GRPO. As shown in Figure 1(e), to introduce controllable stochasticity in the rollout process, SofT-GRPO samples groups of soft-thinking reasoning paths by injecting Gumbel noises into the output token probabilities and employs the Gumbel-Softmax technique to avoid invalid soft-tokens outside the pre-trained discrete-token embedding space. To accurately assign the reward credits to the output probability of LLMs in soft-thinking policy updates, SofT-GRPO leverages the reparameterization trick on the Gumbel distribution. We perform thorough evaluations of SofT-GRPO on three representative LLMs from 1.5B to 7B parameters and across five numerical reasoning benchmarks. Experimental results show that SofT-GRPO-enhanced soft-thinking not only surpasses discrete-token GRPO on the averaged Pass@1 but also yields substantial gains on Pass@32, highlighting the practical advantages of robust policy optimization for soft-thinking reasoning. Our contributions are as follows:
  • We present SofT-GRPO, a powerful RLVR algorithm designed with the soft-thinking reasoning paradigm. It adopts the Gumbel-Softmax technique into the group rollout process, actively obtaining diverse but valid soft-thinking reasoning paths.
  • SofT-GRPO proposes a novel gradient estimation approach via Gumbel reparameterization, first achieving precise attribution of improvements to the LLM’s output probability distributions in policy optimization.
  • Over eight in-domain and out-of-domain benchmarks, SofT-GRPO improves the LLM’s reasoning ability with soft-thinking, outperforming the discrete-token GRPO, especially at higher sample rates (Pass@16 or Pass@32).
2 Preliminaries
2.1 Discrete-Token CoT Reasoning
To solve a $|\boldsymbol{Q}|$-token question $\boldsymbol{Q}=(q_{1},\ldots,q_{|\boldsymbol{Q}|})$, the discrete-token CoT reasoning process generates $|\boldsymbol{R}|$ reasoning CoT tokens $\boldsymbol{R}=(r_{1},\ldots,r_{|\boldsymbol{R}|})$ before predicting the answer $\boldsymbol{A}=(a_{1},\ldots,a_{|\boldsymbol{A}|})$ (Guo et al., 2025; Sprague et al., 2024). Each of $\boldsymbol{Q}$, $\boldsymbol{R}$, and $\boldsymbol{A}$ is a token sequence over the vocabulary $\mathcal{T}$ (i.e., $\boldsymbol{Q},\boldsymbol{R},\boldsymbol{A}\in\mathcal{T}^{*}$), where $\mathcal{T}$ is the set of all possible language tokens. Language reasoning tokens and answer tokens are generated with the next-token prediction (NTP) policy of LLM $\pi_{\theta}$ as follows:
$p(\boldsymbol{R},\boldsymbol{A}|\boldsymbol{Q})=$ $\prod_{t=1}^{|\boldsymbol{R}|}\pi_{\theta}(r_{t}|[\boldsymbol{Q},(r_{1},\ldots,r_{t-1})])$ (1)
$\quad\prod_{t=1}^{|\boldsymbol{A}|}\pi_{\theta}(a_{t}|[\boldsymbol{Q},\boldsymbol{R},(a_{1},\ldots,a_{t-1})]),$
where $\cdot$,\ $\cdot$] and $\cdot$,\ $\cdot$,\ $\cdot$] denotes concatenation.
$\mathcal{J}_{\text{GRPO}}(\theta)=\frac{1}{G}\mathbb{E}_{\boldsymbol{Q}\sim\mathcal{D},\{\boldsymbol{R}\}_{g=1}^{G},\{\boldsymbol{A}\}_{g=1}^{G}\sim p(\cdot,\cdot|\boldsymbol{Q})}$ (2)
$\Bigg[\sum_{g=1}^{G}\frac{1}{\left|\boldsymbol{R}_{g}\right|+\left|\boldsymbol{A}_{g}\right|}\sum_{t=1}^{\left|\boldsymbol{R}_{g}\right|+\left|\boldsymbol{A}_{g}\right|}\Big(\min\left(p_{g,t}\hat{A}_{g},\text{clip}(p_{g,t},1-\epsilon,1+\epsilon)\hat{A}_{g}\right)-\beta\mathbb{D}_{KL}(\pi_{\theta}||\pi_{\theta_{\text{ref}}})\Bigg]$
[…]
Supervised Fine-tuning (SFT) for Discrete-Token Reasoning. As straightforward fine-tuning methods for LLM reasoning, the SFT methods (Wu et al., 2025b; Zheng and Lee, 2025) collect high-quality CoT labels for each question and then fine-tune LLMs for correct predictions. However, SFT methods highly rely on the quality of CoT labels, which are hard to obtain for complex datasets. Moreover, there are also concerns about the out-of-domain generalization ability of SFT (Chu et al., 2025).
RLVR Fine-tuning for Discrete-Token Reasoning. RLVR fine-tuning methods such as GRPO (Liu et al., 2024), Dr. GRPO (Liu et al., 2025b), and DAPO (Yu et al., 2025) sample several CoTs $\boldsymbol{R},\boldsymbol{A}]$ and assign a reward to each of them based on the quality of the answers $\boldsymbol{A}$. The standard discrete-token GRPO algorithm (Shao et al., 2024) samples $G$ CoTs for each question $\boldsymbol{Q}$ with sampling policy $\pi_{\theta_{\text{old}}}$ and optimizes the reasoning policy towards CoTs with higher rewards. Its loss function is shown in Eq. (2), where $\hat{A}_{g}$ represents the advantage function for the $g$-th CoT, the reward function $f(\boldsymbol{A}_{g})=1$ if and only if the answer $\boldsymbol{A}_{g}$ is correct, $\epsilon$ is the clipping hyperparameter, $p_{g,t}$ is calculated in an off-policy way for both the current policy $\pi_{\theta}$ and the sampling policy $\pi_{\theta_{\text{old}}}$, and $\mathbb{D}_{KL}$ represents the KL-divergence of $\pi_{\theta}$ and a reference policy $\pi_{\theta_{\text{ref}}}$. RLVR methods can improve the discrete-token reasoning performance of LLMs across a wide range of reasoning problems without the need for CoT labels. However, for each CoT in each LLM step, discrete-token GRPO updates the probability of only one token based on the reward. As mentioned in Yue et al. (2025), making RLVR does not really incentivize reasoning capacity in LLMs beyond the base model.
2.2 Soft-Thinking Reasoning Paradigm
At each of the CoT reasoning steps, conventional discrete-token reasoning is constrained to selecting one token from the token set $\mathcal{T}$, which may hinder the model’s ability to express certain abstract concepts that cannot be easily represented by a single deterministic token (Zhang et al., 2025b), thus undermining the LLM’s expression ability in reasoning. To represent abstract concepts, Zhang et al. (2025b) presents the soft-thinking paradigm, which replaces discrete-token reasoning steps $\boldsymbol{R}$ with a soft-thinking reasoning path $\boldsymbol{S}=(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{|\boldsymbol{S}|})$. Each token $\boldsymbol{s}_{i}\in\mathcal{R}^{d}$ (noted soft token) is a real vector, which is the sum of token embeddings weighted by their respective output probabilities. They are then fed into the next LLM step as follows:
$\boldsymbol{p}_{t}=\pi_{\theta}(\cdot|[\boldsymbol{Q},(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{t-1})])),$ (3)
$\boldsymbol{s}_{t}=\sum_{i=1}^{|\mathcal{T}|}p_{i}\cdot\boldsymbol{e}_{i},$
where $p_{i}\in[0,1]$ is the predicted probability of token $i$ in $\boldsymbol{p}_{t}$ and $\boldsymbol{e}_{i}\in\mathcal{R}^{d}$ is the LLM embeddings of token $i$. As detailed in Appendix A.3, the soft tokens lie in the convex hull of token embeddings, making it possible to maintain validity while transmitting more information.
Based on Eq. (3), Wu et al. (2025a) find that most pre-trained LLMs tend to be single-threaded and the vanilla soft-thinking reasoning may lead to greedy feedback loop that suppresses alternative reasoning paths and undermines the benefits of transmitting more informative soft-tokens. So, they propose to introduce sampling strategies in the soft-thinking reasoning process for randomness, employing Gumbel noises and Gumbel-Softmax technique (Maddison et al., 2016; Jang et al., 2016) based on the output probability $\boldsymbol{p}_{t}$ as follows:
$g_{i}=\log p_{i}+\epsilon_{i},$ $y_{i}=\dfrac{\exp(g_{i}/\tau_{g})}{\sum_{i=1}^{|\mathcal{T}|}\exp(g_{i}/\tau_{g})},$ (4)
$\boldsymbol{s}_{t}=\sum_{i=1}^{|\mathcal{T}|}y_{i}\cdot\boldsymbol{e}_{i},$
where $\epsilon_{i}$ is a scalar noise sampled from the Gumbel distribution Gumbel(0,1)(0,1), and $\tau_{g}$ is the temperature of Gumbel-Softmax. Besides Gumbel-Softmax, Wu et al. (2025a) also tries to use the Dirichlet resampling technique as follows:
$(x_{1},\ldots,x_{|\mathcal{T}|})\sim\text{Dirichlet}(\alpha,\boldsymbol{p}_{t}),\quad\boldsymbol{s}_{t}=\sum_{i=1}^{|\mathcal{T}|}x_{i}\cdot\boldsymbol{e}_{i},$ (5)
where $\alpha$ is a scaling parameter. Empirically, incorporated with Gumbel noises and the Gumbel-Softmax technique, the soft-thinking pattern can outperform conventional discrete-token CoT on a broad range of tasks, including numerical, code, and scientific reasoning, without requiring any fine-tuning. (Wu et al., 2025a).
Figure 2: The pipeline of the proposed SofT-GRPO algorithm. In training with each Query $\boldsymbol{Q}$, the SofT-GRPO first generates a group of $G$ soft-thinking reasoning paths with Gumbel noises and the Gumbel-Softmax technique (Wu et al., 2025a). We transmit the value $g^{\prime}_{i}$ and $y^{\prime}_{i}$ for the loss calculation afterward. Then, we reconstruct the soft-thinking input. Finally, we update the soft-thinking policy with the off-policy proximal policy optimization (Schulman et al., 2017), optimizing soft-thinking policies with Gumbel reparameterization.
2.3 Related Work on RLVR Soft-Thinking
However, after being boosted with the RLVR fine-tuning, discrete-token reasoning will clearly outperform soft-thinking. So, Butt et al. (2025) tries to similarly improve the soft-thinking reasoning ability using GRPO by adding Gaussian noise on the soft-token $\boldsymbol{s}_{t}$ in Eq. (3) as follows:
$\hat{\boldsymbol{s}}_{t}=\boldsymbol{s}_{t}+\mathcal{N}(0,\sigma^{2}Id).$ (6)
In deriving the soft-thinking probability in GRPO, they restore the value of $\hat{\boldsymbol{s}}_{t}$ in the rollout and calculate the log probability for the soft-thinking reasoning path $\boldsymbol{S}$ with the Gaussian reparameterization trick as follows, keeping the updating process of the answering part $\boldsymbol{A}$ unchanged:
[…]
Although this method can improve the soft-thinking ability of base LLMs and show improvements compared to discrete-token GRPO on the Pass@32 metrics (i.e., the pass rate with 32 attempts), there is a severe degradation in the average accuracy. This is mainly due to the following two drawbacks: (1) Adding noises on soft-tokens may help avoid inputs outside the pre-trained discrete-token embedding space, but adding noise to soft-tokens instead of logits is not direct and may theoretically mismatch the LLM predictions. As analyzed in Appendix C.1, token embeddings are linearly dependent ($|\mathcal{T}|\gg d$), so they cannot attribute which probability $p_{i}$ is beneficial to the effectiveness of added noise. Moreover, the added noise may even be impossible to be represented by the embeddings. (2) They do not use advanced sampling methods for effectiveness in training (e.g., incorporating Dirichlet resampling or Gumbel-Softmax technique (Wu et al., 2025a)), which will undermine the performance of sampled soft-thinking reasoning paths.
As a result, existing attempts of applying GRPO to soft-thinking will not keep its advantage over the discrete-token CoT under the no-finetune setting.
$\mathcal{J}_{\text{SofT-GRPO}}(\theta)=\frac{1}{G}\mathbb{E}_{\boldsymbol{Q}\sim\mathcal{D},\{\boldsymbol{S}\}_{g=1}^{G},\{\boldsymbol{A}\}_{g=1}^{G}\sim p(\cdot,\cdot|\boldsymbol{Q})}$ (8)
$\Bigg[\sum_{g=1}^{G}\frac{1}{\left|\boldsymbol{S}_{g}\right|+\left|\boldsymbol{A}_{g}\right|}\sum_{t=1}^{\left|\boldsymbol{S}_{g}\right|+\left|\boldsymbol{A}_{g}\right|}\Big(\min\left(p_{g,t}\hat{A}_{g},\text{clip}(p_{g,t},1-\epsilon,1+\epsilon)\hat{A}_{g}\right)-\beta\mathbb{D}_{KL}(\pi_{\theta}||\pi_{\theta_{\text{ref}}})\Bigg]$
[…]
3 SofT-GRPO: Reinforcing Soft-Thinking Policy with Gumbel Reparameterization
To fully unlock the potential of soft-thinking for better LLM reasoning, we present the RLVR algorithm SofT-GRPO. As shown in Figure 2, it first samples soft-thinking reasoning paths with controllable randomness using the Gumbel-Softmax technique (Wu et al., 2025a). In the following policy update stage, we propose a novel SofT-GRPO loss function with the Gumbel reparameterization trick, which accurately assigns reward credits to token probabilities.
3.1 Group Rollout with Gumbel Noise
As an off-policy RLVR algorithm, SofT-GRPO samples a group of $G$ soft-thinking CoTs for each query $\boldsymbol{Q}$ during rollout and saves them in a replay buffer. A principal challenge for applying RLVR to soft-thinking is that vanilla soft-thinking generates deterministic soft tokens, as each $\boldsymbol{s}_{t}$ is a fixed weighted sum of embeddings based solely on model probabilities (Eq. (3)). So, to explore diverse and powerful soft-thinking reasoning paths in the rollout process, we introduce the Gumbel-Softmax resampling technique (shown in Eq. (4)) as follows:
$(p_{1},\ldots,p_{|\mathcal{T}|})=\pi_{\theta_{\text{old}}}(\cdot|[\boldsymbol{Q},(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{t-1})]),$ (9)
$g^{\prime}_{i}=\log\ p_{i}+\epsilon_{i},\quad y^{\prime}_{i}=\dfrac{\exp(g^{\prime}_{i}/\tau_{g})}{\sum_{i=1}^{|\mathcal{T}|}\exp(g^{\prime}_{i}/\tau_{g})},$
$\boldsymbol{s}_{t}=\sum_{i=1}^{|\mathcal{T}|}y^{\prime}_{i}\cdot\boldsymbol{e}_{i},$
where $\boldsymbol{e}_{i}$ is the embedding of the $i$-th token, $\pi_{\theta_{\text{old}}}$ is the sampling policy used in rollout, and $\epsilon_{i}$ is a scalar Gumbel noise. Using the inverse transform sampling (Jang et al., 2016), we sample $\epsilon_{i}=-\log(-\log(u))$ where $u\sim\text{Uniform}(0,1)$. Besides bringing stochasticity, as discussed in Section 2.2, resampling with Gumbel-Softmax will bring potential improvements on the quality of reasoning paths.
Due to limitations in the amount of fine-tuning data, fine-tuning LLMs with soft-thinking needs to preserve the pre-trained knowledge. So, applying RLVR to soft-thinking should also keep the generated soft tokens $\boldsymbol{s}_{t}$ within the pre-trained discrete-token embedding space. In contrast to noise injection based on other distributions (e.g., Dirichlet or Gaussian), Gumbel-Softmax can better maintain the stability of the soft-token embeddings when providing stochasticity. As established in Theorem 10 (see Appendix D for proof), it produces a relaxation that maintains samples close to the discrete token distribution, thus mitigating the risk of generating out-of-vocabulary soft tokens.
Theorem 3.1 (Gumbel-max Trick).
Let $(p_{1},\dots,p_{n})$ be nonnegative, and $\epsilon_{1},\dots,\epsilon_{n}$ independent samples from $\mathrm{Gumbel}(0,1)$ (Maddison et al., 2016),
$\Pr\left(j=\arg\max_{i}\left(\log p_{i}+\epsilon_{i}\right)\right)=\frac{p_{j}}{\sum_{i=1}^{n}p_{i}}.$ (10)
3.2 Gumbel Reparameterization for Loss Function
In RLVR with discrete tokens (Eq. (1)), the probability of CoT trajectories can be obtained from the output Categorical distributions. Calculating such a probability is impossible in vanilla soft-thinking. To address this, SofT-GRPO proposes using the Gumbel reparameterization trick after obtaining the Gumbel noise $\boldsymbol{\epsilon}$ during the rollout process. We take $\boldsymbol{\epsilon}$ as RLVR actions, consider transforming $\boldsymbol{\epsilon}$ to soft tokens $\boldsymbol{s}_{t}$ as a part of the transition, and estimate the next-soft-token probability with the density of Gumbel noise as follows:
$p(\boldsymbol{g}^{\prime}|[\boldsymbol{Q},(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{t-1})],\theta_{\text{old}})=\exp\big[\sum_{i=1}^{|\mathcal{T}|}-\epsilon_{i}-\exp(-\epsilon_{i})\big],$ (11)
where $\boldsymbol{g}^{\prime}=(g^{\prime}_{1},\ldots,g^{\prime}_{|\mathcal{T}|})$ is the $g^{\prime}_{i}$ values in Eq. (9) collected from the replay buffer. We provide the derivation for this equation in Appendix C.3. Following GRPO, SofT-GRPO requires off-policy probabilities for clipping. When calculating off-policy probabilities, we first reconstruct the soft-thinking reasoning paths $\boldsymbol{S}=(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{|\boldsymbol{S}|})$ from the restored $y^{\prime}_{i}$ values in the rollout process. Then, the probability is calculated with the restored $g^{\prime}_{i}$ values as follows:
$p(\boldsymbol{g}^{\prime}|[\boldsymbol{Q},(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{t-1})],\theta)=$ (12)
$\qquad\exp\left(\sum_{i=1}^{|\mathcal{T}|}-(g^{\prime}_{i}-\log p_{i})-\exp(-(g^{\prime}_{i}-\log p_{i}))\right),$
$\text{where}\ (p_{1},\ldots,p_{|\mathcal{T}|})=\pi_{\theta}(\cdot|[\boldsymbol{Q},(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{t-1})]).$
Finally, the SofT-GRPO loss function is Eq. (8), where $\pi$ and $\pi_{\text{ref}}$ are token probability distributions that take history soft-tokens $\boldsymbol{S}$ and answer tokens $\boldsymbol{A}$ as input. $p_{g,t}$ is calculated as the probability ratio between current policy $p_{\theta}$ (Eq. (11)) and sampling policy $p_{\theta_{\text{old}}}$ (Eq. (12)). Intuitively, SofT-GRPO recalculates the noise corresponding to $\boldsymbol{g}^{\prime}$ obtained in rollout and encourages this noise only when the current soft-thinking trajectory leads to a higher $\hat{A}_{g}$. Unlike discrete-token GRPO in Eq. (2), SofT-GRPO updates the probability of several tokens at each LLM step $t$.
Table 1: Experiment results of baselines and the proposed SofT-GRPO on five numerical reasoning benchmarks. We cover 3 LLMs from 1.5B to 7B and two reasoning patterns, i.e., discrete-token CoT and soft-thinking reasoning. @1 metrics denote the Mean@32 values, where we run each method 32 times on the dataset for average Pass@1 accuracies. @16 and @32 denote the Pass@16 and Pass@32 values on the dataset, respectively. We multiply all the results by 100 to highlight the differences between results. The best result on each metric and dataset is underlined, the best average result is bolded, and the second-best average result is shaded.
Dataset AIME2024 AIME2025 AMC23 MATH-500 GSM8K Average
Metrics @1 @16 @32 @1 @16 @32 @1 @16 @32 @1 @16 @32 @1 @16 @32 @1 @16 @32
DeepSeek-R1-Distill-Qwen-1.5B Base LLM
Discrete-Token CoT Reasoning Pattern
No-Finetune 30.6 70.0 73.3 23.0 46.7 53.3 70.7 92.5 95.0 84.6 97.8 97.8 81.5 95.8 96.7 58.09 80.54 83.23
+ GRPO 31.8 66.7 76.7 25.3 46.7 46.7 77.3 95.0 95.0 87.1 97.4 97.8 84.9 95.1 95.8 61.28 80.16 82.39
Soft-Thinking Reasoning Pattern
No-Finetune 27.3 66.7 70.0 23.8 46.7 53.3 69.9 95.0 95.0 79.4 93.2 96.6 81.0 94.6 97.1 56.28 79.23 82.41
+ GRPO 29.2 70.0 73.3 25.4 46.7 53.3 75.8 95.0 95.0 86.3 96.8 98.2 84.9 95.6 96.4 60.31 80.81 83.26
+ SofT-GRPO 32.6 76.7 80.0 26.1 50.0 53.3 76.4 97.5 97.5 86.3 97.4 98.0 85.5 96.1 97.0 61.39 83.54 85.18
LLaMA-3.2-3B-Instruct Base LLM
Discrete-Token CoT Reasoning Pattern
No-Finetune 4.4 20.0 26.7 0.3 0.3 1.0 18.3 65.0 75.0 38.1 75.6 84.0 67.9 92.1 94.6 25.79 50.61 56.26
+ GRPO 7.3 23.3 26.7 0.5 3.3 3.3 27.3 62.5 67.5 48.3 77.2 82.6 79.6 95.4 96.5 32.60 52.35 55.32
Soft-Thinking Reasoning Pattern
No-Finetune 3.4 16.7 16.7 0.2 6.7 6.7 17.6 70.0 77.5 36.7 76.0 81.4 66.9 91.6 94.7 24.96 52.18 55.39
+ GRPO 8.0 20.0 23.3 0.7 3.3 10.0 27.3 70.0 75.0 47.8 76.8 81.8 79.2 94.8 96.3 32.60 53.00 57.28
+ SofT-GRPO 7.7 23.3 26.7 0.3 10.0 10.0 31.3 67.5 67.5 47.2 77.6 83.4 77.6 96.4 97.7 32.83 54.96 57.06
DeepSeek-R1-Distill-Qwen-7B Base LLM
Discrete-Token CoT Reasoning Pattern
No-Finetune 55.7 80.0 80.0 39.4 66.7 66.7 89.9 97.5 97.5 93.6 98.8 99.2 89.5 96.7 97.0 73.62 87.94 88.07
+ GRPO 54.0 80.0 80.0 40.4 66.7 70.0 89.2 95.0 95.0 93.5 98.4 99.0 91.4 96.1 96.6 73.69 87.24 88.12
Soft-Thinking Reasoning Pattern
No-Finetune 55.3 80.0 83.3 39.2 66.7 70.0 90.2 95.0 97.5 93.3 98.8 99.0 89.3 96.6 97.0 73.44 87.41 89.38
+ GRPO 55.5 80.0 80.0 37.8 63.3 66.7 90.1 97.5 97.5 93.6 98.6 99.0 91.8 96.7 96.9 73.76 87.23 88.01
+ SofT-GRPO 53.2 80.0 83.3 40.4 60.0 73.3 89.6 97.5 97.5 93.3 98.6 99.0 92.1 97.2 97.7 73.74 86.66 90.18
4 Experiments
In this section, we implement the proposed SofT-GRPO algorithm to reinforce the soft-thinking reasoning of three LLMs, including DeepSeek-R1-Distill-Qwen-1.5B, LLaMA-3.2-3B-Instruct, and DeepSeek-R1-Distill-Qwen-7B.
4.1 Implementation Details
Detailed Settings. In practice, we follow the general setting in soft-thinking works (Wu et al., 2025a), enabling both the top-p and top-k sampling strategies, considering only tokens with $k$ highest probabilities instead of the whole token set $|\mathcal{T}|$ and normalizing their probabilities. In both training and inference, we set top-p as $0.95$ and top-k as $5$, the temperature of LLMs $\tau=1$, and the temperature in Gumbel-Softmax $\tau_{g}=0.1$.
Training & Testing Settings. We employ DeepScaler (Luo et al., 2025) as the training dataset, which contains 40,315 queries. In implementing SofT-GRPO, we use SGLang (Zheng et al., 2024) for the rollout process and the verl-0.4.x framework (Sheng et al., 2024) for RLVR.
We involve five numerical reasoning benchmarks as in-domain test sets (AIME2024, AIME2025, AMC23, MATH-500, and GSM8K). We also employ a scientific reasoning benchmark (GPQA Diamond) and two code benchmarks (HumanEval and MBPP) for out-of-domain evaluation. The maximum generation length is 8192 in training and 32768 in testing. Test answers are verified using the Math Verify package (Kydlíček, 2025). All experiments are implemented on a node of 8× NVIDIA H200 GPUs (141 GB VRAM each), and detailed parameters are shown in Appendix D.
Table 2: Average accuracies on out-of-domain datasets. We cover GPQA Diamond, HumanEval, and MBPP. @1 metrics denote the Mean@32 values, where we run the methods 32 times on the dataset for average Pass@1 accuracies. @8, @16, and @32 denote the Pass@8, Pass@16, and Pass@32 values on the dataset, respectively.
Dataset GPQA Diamond HumanEval MBPP Average
Metrics @1 @8 @16 @32 @1 @8 @16 @32 @1 @8 @16 @32 @1 @8 @16 @32
DeepSeek-R1-Distill-Qwen-1.5B Base LLM
Discrete-Token CoT Reasoning Pattern
No-Finetune 36.7 84.3 92.4 96.0 68.1 87.2 90.2 93.9 65.5 84.8 89.1 91.1 56.77 85.45 90.59 93.64
+ GRPO 35.4 77.8 88.4 93.4 72.2 90.9 92.7 94.5 68.1 85.2 87.2 90.0 58.56 84.62 89.41 92.65
Soft-Thinking Reasoning Pattern
No-Finetune 36.0 83.8 91.9 97.0 67.2 89.6 91.5 92.7 64.7 84.4 86.0 88.7 55.98 85.97 89.79 92.79
+ GRPO 36.5 81.3 91.4 94.4 71.8 89.0 92.7 95.1 68.1 84.8 87.9 90.3 58.82 85.05 90.68 93.28
+ SofT-GRPO 37.3 82.8 89.9 95.5 71.2 88.4 91.5 94.5 68.8 88.7 90.3 91.4 59.08 86.65 90.54 93.80
Table 3: Experiments on using majority voting to boost the performances on AIME2024, AIME2025, AMC23 and GSM8K. @1 represents the Pass@1 result from Table 1, which is averaged over 32 runs. M@16 and M@32 represent Major@16 and Major@32, respectively.
Dataset AIME2024 AIME2025 AMC23 MATH-500 GSM8K Average
Metrics @1 M@16 M@32 @1 M@16 M@32 @1 M@16 M@32 @1 M@16 M@32 @1 M@16 M@32 @1 M@16 M@32
DeepSeek-R1-Distill-Qwen-1.5B Base LLM
Discrete-Token CoT Reasoning Pattern
No Fine-tune 30.6 56.7 60.0 23.0 36.7 40.0 70.7 90.0 95.0 84.6 92.6 92.4 81.5 89.5 89.5 58.1 73.1 75.4
+ GRPO 31.8 53.3 50.0 25.3 36.7 30.0 77.3 92.5 92.5 87.1 91.6 91.4 84.9 90.1 90.4 61.3 72.8 70.9
Soft-Thinking Reasoning Pattern
No Fine-tune 27.3 56.7 60.0 23.8 33.3 33.3 69.9 95.0 92.5 79.4 88.8 89.8 81.0 89.3 89.3 56.3 72.6 73.0
+ GRPO 29.2 43.3 46.7 25.4 30.0 30.0 75.8 92.5 92.5 86.3 91.0 91.4 84.9 90.5 90.8 60.3 69.5 70.3
+ SofT-GRPO 32.6 60.0 63.3 26.1 33.3 36.7 76.4 92.5 95.0 86.3 92.2 92.0 85.5 90.6 90.5 61.4 73.7 75.5
Baselines. We use both conventional discrete-token CoT reasoning and soft-thinking reasoning as baselines for SofT-GRPO. For discrete-token CoT, we include discrete-token GRPO trained on the same dataset and base LLMs without fine-tuning as baselines. For baselines with soft-thinking, we implement the soft-thinking reasoning pattern with the Gumbel-Softmax technique, using the best temperature setting $\tau_{g}=0.5$ proposed in Wu et al. (2025a). Under both patterns, we follow the parameters in (Wu et al., 2025a), setting the temperature $\tau$ of LLMs to 0.6 (we discuss this setting in Appendix E.2), top-p to 0.95, and top-k to 30.
Metrics. We adopt the general metrics of Mean@32 (the average Pass@1 accuracy over 32 runs), Pass@16, and Pass@32, which measure the average probability of covering the correct answer within 16 and 32 runs, respectively.
4.2 Main Result
As shown in Table 1, SofT-GRPO can lead to a clear and consistent improvement from the No-Finetune results under the soft-thinking reasoning pattern. Moreover, SofT-GRPO outperforms applying soft-thinking on top of the discrete-token trained GRPO on average, demonstrating superior soft-thinking reasoning ability.
For Pass@1, SofT-GRPO consistently outperforms the discrete-token trained GRPO across LLMs of three sizes (+0.13% on average accuracy). In contrast, comparing the +GRPO results on the discrete-token and soft-thinking patterns, performing soft-thinking on top of the discrete-token trained GRPO gives inconsistent improvement.
As observed in Yue et al. (2025), GRPO can lead to decreased performance on Pass@K over the No-Finetune model, even though it improves performance for Pass@1. We can observe the same result in Table 1. In contrast, the SofT-GRPO leads to a clear improvement in the Pass@16 (+1.80% on average) and Pass@32 (+2.19% on average) metrics. Appendix E.2 shows that the superiority will also hold for different temperatures. As detailed in Appendix F, the improvement in Pass@16 and Pass@32 is likely due to the fact that for each sample in each LLM reasoning step, SofT-GRPO can reinforce the output probability of several tokens instead of focusing on one in GRPO.
4.3 Comparison in Out-of-Domain Datasets
Besides numerical reasoning, we conduct out-of-domain experiments to evaluate the general reasoning ability of SofT-GRPO. As shown in Table 2, the LLMs fine-tuned with SofT-GRPO on numerical questions can still demonstrate advantages from Pass@1 to Pass@32 on a scientific reasoning benchmark (GPQA Diamond) and two code benchmarks (HumanEval and MBPP).
4.4 Boosting SofT-GRPO with Majority Voting
To further exploit the advantage of SofT-GRPO on Pass@32, in this subsection, we design to boost SofT-GRPO with majority voting (Chen et al., 2024). As shown in Table 3, SofT-GRPO with majority-voting can outperform No-Finetune LLM and LLM fine-tuned with discrete-token GRPO across Major@16 (the accuracy of the most common answer in 16 runs) and Major@32. The results show that SofT-GRPO-fine-tuned LLMs can be strengthened into better reasoning solvers with majority voting.
Table 4: Comparison of the proposed SofT-GRPO to the RLVR fine-tuning method proposed in (Butt et al., 2025). We compare to their reported results under the same training dataset and LLM. Soft Tokens* represents their reported results fine-tuned under the soft-thinking pattern.
Dataset MATH-500 GSM8K
Metrics @1 @32 @1 @32
LLaMA-3.2-3B-Instruct Base LLM
Discrete-Token CoT Reasoning Pattern
No-Finetune 38.1 84.0 67.9 94.6
+ GRPO 48.3 82.6 79.6 96.5
Soft-Thinking Reasoning Pattern
+ Soft Tokens* 41.3 77.9 75.5 95.2
47.2 83.4 77.6 97.7
+ SofT-GRPO (+5.9) (+5.5) (+2.1) (+2.5)
4.5 Comparison to Butt et al. (2025)
We conduct comparison experiments with an existing RLVR algorithm for the soft-thinking pattern, the method in Butt et al. (2025). We refer to this method as Soft Tokens in Table 4, where the proposed SofT-GRPO can demonstrate a clear improvement compared to the results reported in Butt et al. (2025). Compared to SofT-GRPO, the algorithm in Butt et al. (2025) also requires transitioning the $d$-dimensional vector between rollout workers and RLVR workers, making implementation more difficult.
Table 5: Ablation studies on the noise added in the proposed SofT-GRPO. We highlight the performance differences on the ablation variants (i.e., adding Dirichlet noises or Gaussian noises).
Dataset AIME2024 AIME2025 AMC23 MATH-500 GSM8K Average
Metrics @1 @16 @32 @1 @16 @32 @1 @16 @32 @1 @16 @32 @1 @16 @32 @1 @16 @32
DeepSeek-R1-Distill-Qwen-1.5B Base LLM
Soft-Thinking Reasoning Pattern
No-Finetune 27.3 66.7 70.0 23.8 46.7 53.3 69.9 95.0 95.0 79.4 93.2 96.6 81.0 94.6 97.1 56.3 79.2 82.4
+ GRPO 29.2 70.0 73.3 25.4 46.7 53.3 75.8 95.0 95.0 86.3 96.8 98.2 84.9 95.6 96.4 60.3 80.8 83.3
+ SofT-GRPO (Original, Gumbel) 32.6 76.7 80.0 26.1 50.0 53.3 76.4 97.5 97.5 86.3 97.4 98.0 85.5 96.1 97.0 61.4 83.5 85.2
25.4 60.0 70.0 22.9 56.7 63.3 68.4 97.5 97.5 83.6 97.0 97.8 83.6 97.0 97.8 56.8 81.6 85.3
+ SofT-GRPO (Dirichlet Noise) (-7.2) (-16.7) (-10) (-3.2) (+6.7) (+10) (-8) (0) (0) (-2.7) (-0.4) (-0.2) (-1.8) (+0.9) (+0.8) (-4.6) (-1.9) (+0.1)
21.9 60.0 73.0 20.0 40.0 43.3 65.6 97.5 97.5 79.9 95.2 96.6 81.8 94.5 95.5 53.8 77.4 81.2
+ SofT-GRPO (Gaussian Noise) (-10.7) (-16.7) (-7) (-6.1) (-10) (-10) (-10.8) (0) (0) (-6.5) (-2.2) (-1.4) (-3.7) (-1.7) (-1.5) (-7.6) (-6.1) (-4)
(a) Training reward curve of variants on added noises
(b) Validation accuracy curve of variants on added noises
(c) Training reward curve of different hyper-parameters
Figure 3: Smoothed training or validation curves of ablation studies (the dashed background contains the data points). (a) discusses the setting of adding Gumbel noise in SofT-GRPO. (b,c) discusses the setting of top-p=0.95 and the Gumbel-Softmax temperature $\tau_{g}=0.1$.
4.6 Comparison in the Token Efficiency
In addition to accuracy, token efficiency is an important metric for LLMs. As shown in Appendix E.1, SofT-GRPO reduces thinking length compared to No-Finetune, and does not greatly increase tokens compared to discrete-token GRPO. Notably, SofT-GRPO leads to a clear reduction in thinking length for the LLaMA-3.2-3B-Instruct model, highlighting its effectiveness in reducing computational costs.
5 Discussion
5.1 Ablation on the Added Noise
In SofT-GRPO, we employ the Gumbel-Softmax technique for controllable stochasticity. As discussed in Section 2.2, adding Dirichlet noise can be another choice, so we compare the performance of the original SofT-GRPO and its two variants, adding Dirichlet noise or Gaussian noise to probabilities $p_{i}$ in training. The reward curve in training is shown in Figure 3(a), the validation curve is shown in Figure 3(b), and the final performance of variants is shown in Table 5. Compared to adding the original Gumbel noises, LLMs do not gain improvements using Dirichlet noise, and adding the Gaussian noise will cause a poorer initial performance.
5.2 Ablation on Hyper-Parameters
SofT-GRPO sets the top-p as 0.95 and the Gumbel temperature as $\tau_{g}=0.1$. To demonstrate the reason for these settings, in Figure 3(c), we compare the training reward curve of the original SofT-GRPO with two variants (varying top-p to 1.0 or $\tau_{g}$ to 0.25). Results show that both variants will cause a collapse in the training process. As analyzed in Appendix E.3, the two variants will cause a substantial increase in the KL divergence between $\pi_{\theta}$ and $\pi_{\theta_{ref}}$. So, we attribute this kind of collapse to the fact that fine-tuning under the soft-thinking pattern may lead to soft-token inputs outside the pre-trained discrete-token embedding space.
5.3 Visualization: Soft-Thinking after SofT-GRPO
In Appendix H, we provide a visualization of the soft-thinking reasoning path after SofT-GRPO fine-tuning. As shown in Figure 11, high probability paths preserve interpretability, and we also observe new emerged distributions with high mass on a token as well as its antonym (Yeah and No), indicating preservation of multiple reasoning branches.
6 Conclusion
This paper presents a powerful RLVR algorithm, SofT-GRPO, to reinforce LLMs under the soft-thinking reasoning pattern. It integrates controllable stochasticity with Gumbel-Softmax and updates the soft-thinking policy with Gumbel reparameterization. It can demonstrate superior numerical, scientific, and code reasoning ability compared to the conventional discrete-token GRPO on Pass@1, especially Pass@16 and Pass@32. It can also be boosted into a better solver with majority voting. This work demonstrates the advantages of soft-thinking in LLM reasoning. Future works include extending the method for models such as Vision Language Models.
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix Contents
  1. 1.
    Related Work ........................................................................................................................................................................A
    1. (a)
      RLVR on Discrete-Token CoT Reasoning
    2. (b)
      Latent Reasoning
    3. (c)
      The Working Mechanism of Soft-thinking
  2. 2.
    Prompt of Language Reasoning and Latent Generation ........................................................................................................................................................................B
  3. 3.
    Motivation and Theoretical Proof ........................................................................................................................................................................C
    1. (a)
      Motivation: Mismatch in Butt et al. (2025) ........................................................................................................................................................................C.1
    2. (b)
      Proof for Theorem 3.1 ........................................................................................................................................................................D
    3. (c)
      Derivation of Eq. (8), Eq. (11) and Eq. (12) ........................................................................................................................................................................C.3
  4. 4.
    Detailed Parameters for Training & Testing ........................................................................................................................................................................D
  5. 5.
    Supplementary of Experiments ........................................................................................................................................................................E
    1. (a)
      Supplementary of Token Efficiency
    2. (b)
      Experiments under Different Temperatures
    3. (c)
      Supplementary Ablation on Hyper-Parameters
    4. (d)
      Pass@K up to K=1024
    5. (e)
      P-values for Significance on Pass@K
  6. 6.
    Discussion & Analysis on Results ........................................................................................................................................................................F
  7. 7.
    Baselines & Datasets & Licenses ........................................................................................................................................................................G
    1. (a)
      Baselines
    2. (b)
      Datasets
    3. (c)
      Inference Framework
    4. (d)
      Licenses
  8. 8.
    Visualization ........................................................................................................................................................................H
Appendix A Related Work
In this section, we will discuss recent developments on discrete-token RLVR, a similar domain of soft-thinking, the latent reasoning methods, and the working mechanism of soft-thinking.
A.1 RLVR on Discrete-Token CoT Reasoning
Deepseek-R1-zero (Guo et al., 2025) has demonstrated remarkable performance with discrete-token RLVR. And there has been a wide collection of RLVR methods for discrete-token CoT policy optimization, such as GRPO (Shao et al., 2024), Dr. GRPO (Liu et al., 2025b), DAPO (Yu et al., 2025), and Lite PPO (Liu et al., 2025c). To further improve these algorithms towards some specific goals, recent works focus on developing RLVR methods for efficient responses (Feng et al., 2025) or controlling the entropy in CoT generation for better exploration-exploitation balance (Hao et al., 2025).
To generate concise discrete-token CoTs, Arora and Zanette (2025) modifies the reward function by adding penalties on the generation length. Qi et al. (2025a) and Dai et al. (2025) turn to formulate the fine-tuning process for concise language CoT as a multi-objective optimization task with the trade-off between token efficiency and accuracy. They truncate the generation of discrete-token CoT at several thinking budgets and optimize the overall performance over them.
RLVR methods with entropy control try to make a better balance between exploration and exploitation. They try to control the entropy in the training process to stimulate the exploration within the groups of CoTs in GRPO (Wang et al., 2025a).
This paper only considers GRPO as the only RLVR pipeline, because we believe these modifications (e.g., considering entropy or DAPO, which proposes to remove the KL divergence term) should be orthogonal to the two major contributions of SofT-GRPO: Sampling with Gumbel-Softmax and credit assignment with Gumbel reparameterization. In the future, we would like to do a comprehensive investigation to check this.
A.2 Latent Reasoning
Similar to the soft-thinking pattern, latent reasoning methods pass continuous vectors between LLM steps. These methods fully decouple the reasoning process from explicit natural language (which soft-thinking does not do) and perform inference in the hidden space of the model. Generally, latent CoT methods are often diverse from each other and can be mainly divided into token-wise auto-regressive methods and auxiliary strategies (Chen et al., 2025; Zhu et al., 2025). Token-wise auto-regressive methods transform the reasoning process into ’soft thoughts’ with dense latent embeddings (Hao et al., 2024) or specialized tokens (e.g., pause (Goyal et al., 2024; Zelikman et al., 2024)). These methods focus on transferring the original reasoning policy in the language domain to a latent embedding space, including curriculum learning (e.g., Coconut (Hao et al., 2024), LightThinker (Zhang et al., 2025a), SIM-COT (Wei et al., 2025)), self-distillation (e.g., CODI (Shen et al., 2025)), and one-shot compression (e.g., CoLaR (Tan et al., 2025), SynAdapt (Wang et al., 2025b), Latent-SFT (Deng et al., 2025)). Auxiliary strategies (e.g., SoftCoT (Xu et al., 2025a)) generate latent embeddings from an auxiliary module and inject them into the frozen main model (Cheng and Van Durme, 2024; Xu et al., 2025b; Su et al., 2025).
Due to the existing token-wise auto-regressive methods completely treating the language CoTs as the label, these methods can hardly surpass or even reach the performance level of reasoning LLMs. So empirically, these methods can effectively improve the token efficiency compared to language CoT, but there is a clear performance drop. Auxiliary strategies, instead, can effectively boost the performance of the original LLM, sacrificing the running efficiency.
Compared to latent reasoning methods mainly aiming at better efficiency, the proposed SofT-GRPO aims at reinforcing the accuracy of the soft-thinking pattern to surpass discrete-token CoT with GRPO on general reasoning tasks.
A.3 The Working Mechanism of Soft-thinking
As illustrated in (Zhang et al., 2025b), soft-thinking is effective because it transforms the exponential path expansion of reasoning into a tractable and parallelizable process, rigorously approximating full path-summation without sacrificing efficiency or accuracy.
  • Continuous Concept Space: Soft Thinking replaces discrete token choices with “concept tokens,” which are probability-weighted mixtures over all token embeddings. This enables the model to represent more abstract and nuanced concepts beyond individual tokens.
  • Parallel Path Exploration: By retaining the full probability distribution at each reasoning step, Soft Thinking allows the model to implicitly explore multiple reasoning trajectories in parallel, rather than committing early to a single discrete path.
Remarkably, LLMs can directly interpret and process soft tokens without the need for additional fine-tuning. Intuitively, this is because soft tokens are constrained to a well-confined space, specifically the convex hull formed by the model’s embedding vectors.
  • First-Order Linearization for Effectiveness: Moreover, the effectiveness of the soft-thinking paradigm can be understood as a recursive first-order (linear) approximation of the full discrete autoregressive path-sum as follows:
    1. 1.
      For discrete reasoning tokens $\boldsymbol{R}=(r_{1},\ldots,r_{|R|})$, answer $\boldsymbol{A}=(a_{1},\ldots,a_{|A|})$, and LLM policy $\pi_{\theta}$, we have:
      $p(\boldsymbol{A}\mid\boldsymbol{Q})=\sum_{R}\prod_{t=1}^{|\boldsymbol{R}|}\pi_{\theta}(r_{t}\mid[\boldsymbol{Q},(r_{1},...,r_{t-1})])\prod_{t=1}^{|\boldsymbol{A}|}\pi_{\theta}(a_{t}\mid[\boldsymbol{Q},\boldsymbol{R},(a_{1},...,a_{t-1})]).$ (13)
    2. 2.
      Soft-Thinking Tokens: At each reasoning step, define the soft token as the expected embedding:
      $p_{t}\sim\pi_{\theta}(\cdot\mid[\boldsymbol{Q},(\boldsymbol{s}_{1},...,\boldsymbol{s}_{t-1})]),\quad\boldsymbol{s}_{t}=\sum_{i=1}^{|\mathcal{T}|}p_{t,i}\cdot e_{i},$ (14)
      where $p_{t}$ is the token probability distribution, $e_{i}$ is the embedding of token $i$.
    3. 3.
      Then we can recursively replace each sum over discrete tokens by soft tokens. For example, for the first reasoning step:
      $p(\boldsymbol{A}|\boldsymbol{Q})$ $=\sum_{r_{1}}\pi_{\theta}(r_{1}|\boldsymbol{Q})\,p(\boldsymbol{A}|\boldsymbol{Q},r_{1})$ (15)
      $\approx p(\boldsymbol{A}|\boldsymbol{Q},\,\mathbb{E}_{r_{1}}[e_{r_{1}}])$ (16)
      $=p(\boldsymbol{A}|\boldsymbol{Q},\,\boldsymbol{s}_{1}).$ (17)
      Here, $\boldsymbol{s}_{1}=\sum_{i=1}^{|\mathcal{T}|}\pi_{\theta}(r_{1}=i|\boldsymbol{Q})\cdot e_{i}$ is the soft token for the first step.
      For later steps, recursively apply the same procedure:
      $p(\boldsymbol{A}|\boldsymbol{Q},\boldsymbol{s}_{1})$ $=\sum_{r_{2}}\pi_{\theta}(r_{2}|\boldsymbol{Q},\boldsymbol{s}_{1})\,p(\boldsymbol{A}|\boldsymbol{Q},\boldsymbol{s}_{1},r_{2})$ (18)
      $\approx p(\boldsymbol{A}|\boldsymbol{Q},\boldsymbol{s}_{1},\boldsymbol{s}_{2}),$ (19)
      $\text{where}\ \boldsymbol{s}_{2}$ $=\sum_{i=1}^{|\mathcal{T}|}\pi_{\theta}(r_{2}=i|\boldsymbol{Q},\boldsymbol{s}_{1})\cdot e_{i}.$ (20)
    4. 4.
      Recursive Soft-Token Approximation: By recursively applying this linearization at each step, we approximate the original intractable path sum by a deterministic soft-token chain:
      $p(\boldsymbol{A}\mid\boldsymbol{Q})\approx\prod_{t=1}^{|\boldsymbol{S}|}\pi_{\theta}(\boldsymbol{s}_{t}\mid[\boldsymbol{Q},(\boldsymbol{s}_{1},...,\boldsymbol{s}_{t-1})])\prod_{t=1}^{|\boldsymbol{A}|}\pi_{\theta}(a_{t}\mid[\boldsymbol{Q},\boldsymbol{S},(a_{1},...,a_{t-1})]).$ (21)
      where $\boldsymbol{S}=(\boldsymbol{s}_{1},...,\boldsymbol{s}_{|\boldsymbol{S}|})$. Thus, each $\boldsymbol{s}_{t}$ replaces the sum over all possible $r_{t}$, acting as a linear surrogate.
However, it is important to note that not all points in the embedding space (i.e., convex hull of token embeddings) are equally interpretable or accessible to the LLM. For particularly smooth or spread-out distributions (making the approximation in Eq. (15) fail), the expressiveness of soft tokens remains limited. We believe this limitation is closely related to the “collapse” phenomenon discussed in Section 5.2, where the model’s reasoning may degrade or repeat when encountering out-of-distribution soft tokens.
Appendix B Prompt of Language Reasoning and Latent Generation
In this part, we show the prompt we adopt for reasoning problems, including the in-domain numerical reasoning, out-of-domain GPQA reasoning, and code reasoning. We inherit the prompts in Zhang et al. (2025b) for out-of-domain benchmarks. We use the same prompt for both the soft-thinking reasoning pattern and the discrete-token reasoning pattern.
[…]
Prompt for GPQA Reasoning. user Please solve the following multiple-choice question. Please show your choice in the answer field with only the choice letter, e.g.,”answer”: ”C”. {Question}
Prompt for Code Reasoning (HumanEval). user Please solve the programming task below in Python. Code should be wrapped in a markdown code block. “‘python {Question} “‘
Prompt for Code Reasoning (MBPP). user Please solve the programming task with test cases below in Python. Make sure your code satisfies the following requirements: 1. The function name and signature must match exactly as specified in the test cases. 2. Your code should be wrapped in a markdown code block without including any test cases. Task: {Question} Test Cases: “‘python {TestCases} “‘
The blue part represents the specific question (query $\boldsymbol{Q}$), the brown part represents the possible test cases provided in the MBPP code reasoning benchmark. We use the code provided in Zhang et al. (2025b) in the verification process of the responses.
Appendix C Motivation and Theoretical Proof
C.1 Motivation: Mismatch in Butt et al. (2025)
Butt et al. (2025) adds Gaussian noise on the soft-token inputs as follows:
$\boldsymbol{s}_{t}=\sum_{i=1}^{|\mathcal{T}|}p_{i}\cdot\boldsymbol{e}_{i},\quad\hat{\boldsymbol{s}}_{t}=\boldsymbol{s}_{t}+\mathcal{N}(0,\sigma^{2}Id).$ (22)
and calculates the log probability as follows:
[…]
Let $\boldsymbol{E}\in\mathbb{R}^{|\mathcal{T}|\times d}$ be the embedding matrix, and each token probability vector $\boldsymbol{p}\in\Delta^{|\mathcal{T}|-1}$ corresponds to the soft input $\boldsymbol{s}=\boldsymbol{E}^{\top}\boldsymbol{p}\in\mathbb{R}^{d}$.
Assume the observed noisy soft input is $\hat{\boldsymbol{s}}$, and we define a likelihood
$\log p(\hat{\boldsymbol{s}}\mid\boldsymbol{p})\propto-\frac{1}{2\sigma^{2}}\|\hat{\boldsymbol{s}}-\boldsymbol{E}^{\top}\boldsymbol{p}\|_{2}^{2}.$ (24)
Suppose we want to regard this as a likelihood on $\boldsymbol{p}$. In general, the mapping $\boldsymbol{p}\mapsto\boldsymbol{E}^{\top}\boldsymbol{p}$ is many-to-one: since $|\mathcal{T}|>d$, the kernel of $\boldsymbol{E}^{\top}$ is nontrivial, so
$\exists\,\boldsymbol{p}_{1}\neq\boldsymbol{p}_{2},\quad\text{with}\quad\boldsymbol{E}^{\top}\boldsymbol{p}_{1}=\boldsymbol{E}^{\top}\boldsymbol{p}_{2}.$ (25)
Thus, the same $\boldsymbol{s}$ may correspond to infinitely many $\boldsymbol{p}$. This means that, under this Gaussian model, two different token mixtures can lead to the exact same log-probability value. The information about the original token distribution is partially lost in the embedding projection, unless $\boldsymbol{E}$ is invertible (which it is not). Thus, the use of a Gaussian noise model on the embedding space gives a mismatch to the true simplex-based probability geometry.
In summary of Drawback 1: Due to the non-injectivity (non-invertibility) of the embedding transformation, the model $\log p(\hat{\boldsymbol{s}}_{t})\propto-\|\hat{\boldsymbol{s}}_{t}-\boldsymbol{s}_{t}\|^{2}$ does not define a true likelihood on the simplex of token probabilities.
The above mismatch is not only due to the non-invertibility of the embedding matrix. Even if we restrict $\boldsymbol{p}$ to be sparse (nonzero only on a top-$k$ set, which is a general setting of LLMs or a common nature of LLM predictions), and even if the corresponding submatrix $\boldsymbol{E}_{\mathcal{K}}$ is invertible, the process of adding Gaussian noise in the $d$-dimensional embedding space fundamentally breaks the connection to sparse token distributions.
More specifically, after adding Gaussian noise, with a general top-k setting (e.g., kk=10 to 30, and $k<<d$) the perturbed embedding $\hat{\boldsymbol{s}}_{t}=\boldsymbol{s}_{t}+\boldsymbol{\epsilon}$ (with $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\sigma^{2}\boldsymbol{I}_{d})$) will almost surely not lie in the convex hull of any set of $k$ token embeddings. In other words,
$\forall~\hat{\boldsymbol{s}}_{t},~\text{for almost all }\boldsymbol{\epsilon},~\nexists~k\text{-sparse }\boldsymbol{p}\text{ such that }\hat{\boldsymbol{s}}_{t}=\boldsymbol{E}^{\top}\boldsymbol{p}.$ (26)
The set of all top-$k$ soft-token embeddings forms a low-dimensional union of simplices in $\mathbb{R}^{d}$, which is a measure-zero subset of the space. The probability of a randomly perturbed embedding $\hat{\boldsymbol{s}}_{t}$ coinciding with a legal top-$k$ mixture is thus zero.
Therefore, defining the likelihood $p(\hat{\boldsymbol{s}}_{t})$ as a simple Gaussian on embedding space cannot be interpreted as a likelihood on the space of top-$k$ soft tokens—not only due to non-invertibility or nonlinearity, but more fundamentally because most $\hat{\boldsymbol{s}}_{t}$ produced by noise are not realizable by any top-$k$ soft-token distribution.
In summary of Drawback 2: Under the general top-k setting, the likelihood $p(\hat{\boldsymbol{s}}_{t})$ as a simple Gaussian on the embedding space cannot be interpreted as a likelihood on the space of top-$k$ embeddings.
C.2 Proof for Theorem 3.1
Theorem 10 (Gumbel-max Trick) Let $(p_{1},\dots,p_{n})$ be nonnegative real numbers, not all zero. Let $g_{1},\dots,g_{n}$ be independent samples from $\mathrm{Gumbel}(0,1)$. Then,
$\Pr\left(j=\arg\max_{i}(g_{i}+\log p_{i})\right)=\frac{p_{j}}{\sum_{i=1}^{n}p_{i}}.$ (27)
Proof.
For any $j\in\{1,\dots,n\}$,
$\Pr\left(j=\arg\max_{i}(g_{i}+\log p_{i})\right)$ $=\Pr\left(g_{j}+\log p_{j}\geq g_{i}+\log p_{i},\,\,\forall i\neq j\right)$
$=\int_{-\infty}^{+\infty}\prod_{i\neq j}\Pr\left(g_{i}\leq g_{j}+\log p_{j}-\log p_{i}\right)f_{g_{j}}(g_{j})\,dg_{j},$
where $f_{g_{j}}(g)=e^{-g-\exp(-g)}$ is the PDF of the standard Gumbel distribution.
$\Pr(g_{i}\leq t)=F_{\mathrm{Gumbel}}(t)=\exp\left(-e^{-t}\right)$, so
$\prod_{i\neq j}\Pr\left(g_{i}\leq g_{j}+\log p_{j}-\log p_{i}\right)$
$=\prod_{i\neq j}\exp\left(-e^{-(g_{j}+\log p_{j}-\log p_{i})}\right)$
$=\exp\left(-\sum_{i\neq j}e^{-(g_{j}+\log p_{j}-\log p_{i})}\right)$
$=\exp\left(-e^{-g_{j}}\sum_{i\neq j}\frac{p_{i}}{p_{j}}\right).$
The total probability is
$\int_{-\infty}^{+\infty}e^{-g_{j}}e^{-e^{-g_{j}}}\exp\left(-e^{-g_{j}}\sum_{i\neq j}\frac{p_{i}}{p_{j}}\right)dg_{j}$
$=\int_{-\infty}^{+\infty}e^{-g_{j}}\exp\left(-e^{-g_{j}}\left[1+\sum_{i\neq j}\frac{p_{i}}{p_{j}}\right]\right)dg_{j}.$
Let $S=1+\sum_{i\neq j}\frac{p_{i}}{p_{j}}=\frac{\sum_{i=1}^{n}p_{i}}{p_{j}}$. Substitute $y=e^{-g_{j}},\,dg_{j}=-\frac{dy}{y},\,y\in(0,+\infty)$,
$=\int_{y=+\infty}^{0}y\exp(-yS)\left(-\frac{dy}{y}\right)=\int_{0}^{+\infty}\exp(-yS)dy=\frac{1}{S}=\frac{p_{j}}{\sum_{i=1}^{n}p_{i}}.$
Thus, combining all the above equations,
$\Pr\left(j=\arg\max_{i}(g_{i}+\log p_{i})\right)=\frac{p_{j}}{\sum_{i=1}^{n}p_{i}}.$
C.3 Derivation of Eq. (8), Eq. (11) and Eq. (12)
The Gumbel-Softmax rollout of SofT-GRPO is given in Eq. (9). On the determinism of the mapping. It is crucial to note that, for fixed logits (or probabilities) $\mathbf{p}_{t}$ and Gumbel noises $\boldsymbol{\epsilon}_{t}$, the resulting soft token $\mathbf{s}_{t}$ is a deterministic function of $(\mathbf{p}_{t},\boldsymbol{\epsilon}_{t})$:
$\mathbf{s}_{t}=\mathrm{SoftEmbed}(\mathbf{p}_{t},\boldsymbol{\epsilon}_{t}),$ (28)
Hence, in our analysis, we always track the generative randomness via $\boldsymbol{\epsilon}_{t}$ (consider $\boldsymbol{\epsilon}_{t}$ as the action in RLVR), and not via $\mathbf{s}_{t}$ directly. Cause every $\boldsymbol{g}^{\prime}$ can map to one deterministic soft-token $\boldsymbol{s}_{t}$, but $\boldsymbol{s}_{t}$ can map to multiple $\boldsymbol{g}^{\prime}$, calculating density on $\boldsymbol{g}^{\prime}$ avoids the ill-posedness of attempting to define sampling densities directly in the soft-token space. So, we consider outputting $\boldsymbol{g}^{\prime}$ as an action and put the generation of $\boldsymbol{s}_{t}$ as a part of the environment transition.
Eq. (11): log-likelihood under the old policy.
During rollout (generated by the old policy $\theta_{\mathrm{old}}$), we sample Gumbel noises $\boldsymbol{\epsilon}_{t}=(\epsilon_{t,1},\dots,\epsilon_{t,|\mathcal{T}|})$. The randomness of $\mathbf{s}_{t}$ is entirely attributable to $\boldsymbol{\epsilon}_{t}$, given $\mathbf{p}_{t}$. The joint density factorizes as
$p(\boldsymbol{\epsilon}_{t})=\prod_{i=1}^{|\mathcal{T}|}f_{\mathrm{Gumbel}}(\epsilon_{t,i}),\qquad f_{\mathrm{Gumbel}}(x)=\exp\!\big(-x-\exp(-x)\big),$
where $f_{\mathrm{Gumbel}}$ represents the probability density function of a Gumbel distribution. Hence, the corresponding log-likelihood is
$\log p(\boldsymbol{\epsilon}_{t})=\sum_{i=1}^{|\mathcal{T}|}\log f_{\mathrm{Gumbel}}(\epsilon_{t,i})=\sum_{i=1}^{|\mathcal{T}|}\Big[-\epsilon_{t,i}-\exp(-\epsilon_{t,i})\Big],$
which gives Eq. (11). In this formulation, the density is over $\boldsymbol{\epsilon}_{t}$, while $\mathbf{s}_{t}$ is merely a deterministic function thereof.
Eq. (12): log-likelihood under the current policy.
For off-policy correction, we reuse the same realized perturbations from rollout. Equivalently, we store the perturbed logits.
$g^{\prime}_{t,i}\triangleq\log p^{\mathrm{old}}_{t,i}+\epsilon_{t,i},$
so that, under a new policy $\theta$ with probabilities $p_{t,i}=\pi_{\theta}(i\mid[Q,(\mathbf{s}_{1},\dots,\mathbf{s}_{t-1})])$, the implied noise that would have generated the same $g^{\prime}_{t,i}$ under the new policy is
$\epsilon_{t,i}=g^{\prime}_{t,i}-\log p_{t,i}.$
Therefore, the log-density of the same realized perturbation under the new policy is
$\sum_{i=1}^{|\mathcal{T}|}\log f_{\mathrm{Gumbel}}(g^{\prime}_{t,i}-\log p_{t,i})=\sum_{i=1}^{|\mathcal{T}|}\Big[-(g^{\prime}_{t,i}-\log p_{t,i})-\exp\!\big(-(g^{\prime}_{t,i}-\log p_{t,i})\big)\Big],$
which is exactly Eq. (12). This preserves consistency in credit assignment and importance sampling, as all stochasticity and probability mass are correctly attributed to $\boldsymbol{\epsilon}_{t}$ rather than $\mathbf{s}_{t}$.
Eq. (8): Off-Policy Ratio
Cause the transformation from $\boldsymbol{g}^{\prime}$ to $\boldsymbol{s}_{t}$ is fixed, the reparameterization process of SofT-GRPO considers the added noise as action. So, using the Eq. (11) and Eq. (12), we can get the final result in Eq. 8 for the soft-thinking part.
$\frac{p(\boldsymbol{g}^{\prime}|[\boldsymbol{Q},(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{t-1})],\theta)}{p(\boldsymbol{g}^{\prime}|[\boldsymbol{Q},(\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{t-1})],\theta_{\text{old}})}=$ $\frac{\exp\left(\sum_{i=1}^{|\mathcal{T}|}-(g^{\prime}_{i}-\log p_{i})-\exp(-g^{\prime}_{i}+\log p_{i})\right)}{\exp\left(\sum_{i=1}^{|\mathcal{T}|}-\epsilon_{i}-\exp(-\epsilon_{i})\right)}$ (29)
$=$ $\exp\left(\sum_{i=1}^{|\mathcal{T}|}\big(-(g^{\prime}_{i}-\log p_{i})-\exp(-g^{\prime}_{i}+\log p_{i})\big)-\big(-\epsilon_{i}-\exp(-\epsilon_{i})\big)\right).$
Appendix D Detailed Parameters for Training & Testing
All our experiments are performed on 8 $\times$ H200 GPUs in about 2 to 3 days. Compared to discrete-token GRPO, SofT-GRPO requires $\times$2 to $\times$3 in time, mainly due to the more time consumption for soft-thinking rollout (Zhang et al., 2025b; Wu et al., 2025b). We show the experimental configurations in 6 for No-Finetune, 7 for GRPO, and 8 for SofT-GRPO.
Table 6: Parameters of No-Finetune
Parameter Value
No-Finetune Testing under Discrete-Token Reasoning Paradigm
Maximum response length $32768$ tokens
Sampling temperature 0.6
(top-p, top-k) (0.95, 30)
No-Finetune Testing under Soft-Thinking Reasoning Paradigm
Maximum response length $32768$ tokens
Sampling temperature 0.6
Gumbel-temperature $\tau_{g}$ 0.5
(top-p, top-k) (0.95, 30)
Table 7: Parameters of Discrete-Token GRPO
Parameter Value
Discrete-Token GRPO Training
Maximum response length $8192$ tokens
Sampling temperature 1.0
(top-p, top-k) (1, -1)
Group Size $G$ 8
Learning rate $1\times 10^{-6}$
KL loss coefficient $\beta$ 0.001
Policy clipping parameter $\epsilon$ 0.2
Discrete-Token GRPO Testing under Discrete-Token Reasoning Paradigm
Maximum response length $32768$ tokens
Sampling temperature 0.6
(top-p, top-k) (0.95, 30)
GRPO Testing under Soft-Thinking Reasoning Paradigm
Maximum response length $32768$ tokens
Sampling temperature 0.6
Gumbel-temperature $\tau_{g}$ 0.5
(top-p, top-k) (0.95, 30)
Table 8: Parameters of SofT-GRPO
Parameter Value
SofT-GRPO Training
Maximum response length $8192$ tokens
Sampling temperature 1.0
(top-p, top-k) (0.95, 5)
Group Size $G$ 8
Gumbel-temperature $\tau_{g}$ 0.1
Learning rate $1\times 10^{-6}$
KL loss coefficient $\beta$ 0.001
Policy clipping parameter $\epsilon$ 0.2
SofT-GRPO Testing under Soft-Thinking Reasoning Paradigm
Maximum response length $32768$ tokens
Sampling temperature 1.0
Gumbel-temperature $\tau_{g}$ 0.1
(top-p, top-k) (0.95, 5)
Appendix E Supplementary of Experiments
Table 9: Experiments on the token efficiency for baselines and the proposed SofT-GRPO. #Token values in Table represent the number of tokens across all queries, and the #Token_c values represent the number of tokens across correct queries.
Dataset AIME2024 AIME2025 AMC23 MATH-500 GSM8K Average
Metrics #Token #Token_c #Token #Token_c #Token #Token_c #Token #Token_c #Token #Token_c #Token #Token_c
DeepSeek-R1-Distill-Qwen-1.5B Base LLM
Discrete-Token CoT Reasoning Pattern
No-Finetune 16241.6 14997.8 16416.3 13448.8 10052.2 9394.2 5616.6 5368.1 1839.3 1772.5 10033.2 8996.3
+ GRPO 8927.6 8535.6 8039.4 6414.2 5001.3 4672.8 3106.1 2958.5 1417.1 1370.6 5298.3 4790.3
Soft-Thinking Reasoning Pattern
No-Finetune 17857.8 16191.3 17569.4 14582.4 11269.9 10482.0 4015.9 3888.3 1699.3 1649.9 10482.5 9358.8
+ GRPO 9383.2 8934.5 8007.2 7325.8 5203.7 4894.4 3233.6 3131.7 1385.5 1349.9 5442.7 5127.2
+ SofT-GRPO 11039.6 10756.1 10519.6 7831.3 5900.2 5630.4 3549.5 3399.2 1577.6 1542.5 6517.3 5831.9
LLaMA-3.2-3B-Instruct Base LLM
Discrete-Token CoT Reasoning Pattern
No-Finetune 5010.5 3334.1 4297.6 2430.9 2790.7 2398.1 1850.1 1494.2 236.6 224.6 2837.1 1976.4
+ GRPO 8991.8 6368.9 8443.8 11088.3 4998.1 3978.5 3334.8 2463.0 502.0 453.6 5254.1 4870.5
Soft-Thinking Reasoning Pattern
No-Finetune 4859.2 2727.7 4934.8 3735.2 3259.9 2624.0 2086.1 1494.1 243.5 227.8 3076.7 2161.8
+ GRPO 9749.0 7627.8 8498.9 10041.4 5389.9 4987.3 3701.9 2764.6 541.9 505.8 5576.3 5185.4
+ SofT-GRPO 829.7 862.9 893.1 879.6 911.5 927.9 632.8 575.7 294.6 292.9 712.3 707.8
DeepSeek-R1-Distill-Qwen-7B Base LLM
Discrete-Token CoT Reasoning Pattern
No-Finetune 13120.7 11511.4 14347.7 11750.7 6346.0 5987.1 3998.9 3939.0 1061.1 1032.0 7774.9 6844.0
+ GRPO 7795.9 7116.1 8003.5 7369.8 4050.9 3719.7 2473.3 2405.5 1146.7 1118.7 4694.0 4346.0
Soft-Thinking Reasoning Pattern
No-Finetune 13017.4 11888.6 14116.8 11507.4 6346.7 6043.6 3947.4 3858.6 996.3 962.2 7684.9 6852.1
+ GRPO 7464.9 6531.4 8291.6 7160.4 3931.7 3736.7 2473.2 2423.0 1117.0 1104.5 4655.7 4191.2
+ SofT-GRPO 8035.8 7556.8 8381.7 7873.7 4008.0 3843.5 2630.2 2583.5 1293.2 1276.4 4869.8 4626.8
E.1 Supplementary of Token Efficiency
As shown in Section 4.6, besides the performance, the token efficiency of LLMs is also an important metric. In this section, we compare the token efficiency of baselines and SofT-GRPO in Table 9. Compared to No-Finetune variants, SofT-GRPO can demonstrate a clear refinement in both the token efficiency across all queries and the token efficiency across correct queries. Compared to the discrete-token GRPO, SofT-GRPO will not cause severe token improvement. Specifically, we observe a severe reduction in the thinking length of the LLaMA-3.2-3B-Instruct model. As shown in Figure 4, unlike GRPO, SofT-GRPO maintains and even enhances the token efficiency compared to the base LLM when training progresses, demonstrating the effectiveness of SofT-GRPO in reducing computational consumption.
Figure 4: Token consumption curve on LLaMA-3.2-3B-Instruct Base LLM during training.
E.2 Experiments under Different Temperatures
In Section 4, we adopt the setting of $\tau=0.6$ for discrete-token CoT. To investigate whether other temperature settings will break our observations in our main experiments, similar to Yue et al. (2025), we try temperatures from the collection of $\tau=0.6$, $\tau=0.8$, $\tau=1.0$, $\tau=1.2$, and $\tau=1.4$.
As shown in Figure 5, we conduct experiments on the 1.5B LLMs (i.e., DeepSeek-R1-Distill-Qwen-1.5B Base LLM) for their average Pass@K accuracies across five numerical reasoning benchmarks (i.e., AIME2024, AIME2025, AMC23, MATH-500, GSM8K), where the proposed SofT-GRPO can demonstrate outstanding results compared to GRPO and No-Finetune variants with various temperatures, from Pass@1 to Pass@32.
Figure 5: Running discrete-token CoT methods (GRPO and No-finetune) with more temperature options on DeepSeek-R1-Distill-Qwen-1.5B Base LLM. Pass@K represents the pass rate within at most k runs, and Pass@1 is additionally averaged from 32 runs. Experiments are run on the five datasets in Table 1 for the average.
E.3 Supplementary Ablation on Hyper-Parameters
(a) Training KL Divergence curve between $\pi_{\theta_{\text{ref}}}$ and $\pi_{\theta}$
(b) Training Proximal Policy Optimization (PPO) (Schulman et al., 2017) KL Divergence curve between $\pi_{\theta_{\text{old}}}$ and $\pi_{\theta}$
Figure 6: KL Divergence curves in SofT-GRPO over DeepSeek-R1-Distill-Qwen-1.5B Base LLM.
In this subsection, we further investigate the hyperparameter settings. As briefly shown in Section 5.2, adopting a higher top-p or $\tau_{g}$ will cause collapses in training. We attribute these collapses to the case that some soft-thinking tokens may become incomprehensible for LLMs.
As shown in Figure 6(a), the variants (varying top-p to 1.0 or varying $\tau_{g}$ to 0.25) will demonstrate higher divergence between the fixed pre-trained $\pi_{\theta_{\text{ref}}}$ and the current policy, which can be an indicator of the inputs outside the pre-trained embedding space.
Recently, Liu et al. (2025a); Qi et al. (2025b) provides excellent insight into the collapse situation in the RLVR fine-tuning. When collapse is caused by a precision issue, they observe a very high KL divergence between $\pi_{\theta_{\text{old}}}$ and $\pi_{\theta}$ (Refer to Figure 3 in (Qi et al., 2025b), $10^{-3}$ even higher). However, in the variants shown in Figure 6(b), we find that their KL divergence in PPO policies is less than $10^{-5}$, indicating that the variants of SofT-GRPO are less likely to have precision issues.
E.4 Pass@K up to K=1024
Table 1 and Figure 5 show the Pass@K for K at most 64. In Yue et al. (2025), they conduct experiments for K up to 1024. We adopt this setting. Confined by limited computational resources, we implement the proposed SofT-GRPO over DeepSeek-R1-Distill-Qwen-1.5B for 1024 runs, as well as the GRPO and No-Finetune baselines with the discrete-token CoT reasoning pattern.
As shown in Figure 7, we calculate Pass@16 to Pass@1024 on AIME2024 and AMC 23. As a difference from the result in Table 1, we obtain Pass@K for each K runs instead of the first K runs. For example, when calculating Pass@32, we divide the 1024 runs into 32 groups, calculating their Pass@K value and averaging. In Table 1 instead, we calculate the Pass@K at first 32 runs.
Figure 7 obtains the same result pattern compared to the results in (Yue et al., 2025), where GRPO will outperform the No-finetune version when K is small, but will be left behind when K increases. SofT-GRPO can consistently outperform the GRPO and no-finetune counterparts on all K (from 16 to 1024).
(a) Pass@16 to Pass@1024 on AIME2024
(b) Pass@16 to Pass@1024 on AMC23
Figure 7: Pass@16 to Pass@1024 on AIME2024 and AMC 23. The fluctuation in Figure (a) occurs because AIME only has 30 instances, and SofT-GRPO correctly solved three out of four (1024/256) problems when K=256. In Figure (b), all algorithms ultimately achieved complete correctness. Consistent with this, we checked CoT to prove that the correct results did not come from random guessing.
E.5 P-values for Significance on Pass@K
Besides the results in Table 1 and Figure 5, in this section, we use results from 1024 runs to demonstrate the significance test for Pass@16 and Pass@32 (dividing 64 groups for Pass@16 from the 1024 runs and 32 groups of Pass@32). We calculated the t-test for the one-sided hypothesis test (See Figure 7, the average result of SofT-GRPO outperforms the GRPO and No-Finetune baselines), and taking 0.05 as the threshold, the results in Figure 8 and 9 show that in Pass@16 with 64 runs, and Pass@32 with 32 runs:
  • SofT-GRPO can significantly outperform GRPO on AIME2024 with Pass@16 (0.0067), AMC23 with Pass@32 (0.022), with quite small values (0.13) on both AIME2024 with Pass@32 and AMC23 with Pass@16.
  • SofT-GRPO can get quite small p-values (0.081 and 0.054) on AMC23 with Pass@16 and Pass@32.
(a) P-scores of Pass@16 over SofT-GRPO, GRPO, and No-finetune on AIME2024
(b) P-scores of Pass@32 over SofT-GRPO, GRPO, and No-finetune on AIME2024
Figure 8: Significance test on AIME2024.
(a) P-scores of Pass@16 over SofT-GRPO, GRPO, and No-finetune on AMC23
(b) P-scores of Pass@32 over SofT-GRPO, GRPO, and No-finetune on AMC23
Figure 9: Significance test on AMC23.
Appendix F Discussion & Analysis on Results
In this section, we will provide a detailed analysis of the improvement in Pass@K. Yue et al. (2025) first validates that RLVR techniques in discrete-token GRPO will not incentivize reasoning capacity in LLMs beyond the base model, demonstrating equal or even inferior Pass@K values when k improves. So, Recent studies (Walder and Karkhanis, 2025; Mahdavi et al., 2025; Peng et al., 2025) try to develop RL methods that optimize the Pass@K accuracy directly. Among them, Walder and Karkhanis (2025); Chow et al. (2024); Mahdavi et al. (2025); Hu et al. (2026) focus on reallocating the reward within the group, encouraging rewards of negative discrete-token CoTs in groups with high Pass@K value. Peng et al. (2025) try to improve the model’s Pass@K performance by avoiding the over-concentration on the top-1 selection within the token set.
In SimKO, at high-entropy (forking) tokens, positive (correct) samples are rewarded not only on the sampled token but spread across the top-K most probable tokens, promoting output diversity and better Pass@K. For negative (incorrect) samples, SimKO applies stronger penalties to the top-1 token and lighter penalties to others, which prevents the output probabilities from collapsing onto a single choice. This entropy-aware and asymmetric strategy encourages exploration, outperforming traditional RL methods that only focus on the sampled token.
In the proposed SofT-GRPO, we have the policy update in the soft-thinking part ($t$-th token) as:
$\nabla_{\theta}L(\theta)$ $=\nabla_{\theta}\left(\sum_{i=1}^{|\mathcal{T}|}\left[-g^{\prime}_{i}+\log p_{i}-\exp(-g^{\prime}_{i}+\log p_{i})\right]-\left[-\epsilon_{i}-\exp(-\epsilon_{i})\right]\right)$
$=\sum_{i=1}^{|\mathcal{T}|}\nabla\theta\left(\log p_{i}-\exp(-g^{\prime}_{i}+\log p_{i})\right)$
$=\sum_{i=1}^{|\mathcal{T}|}\left(\nabla_{\theta}\log p_{i}-\exp(-g^{\prime}_{i}+\log p_{i})\nabla_{\theta}\log p_{i}\right)$
$=\sum_{i=1}^{|\mathcal{T}|}\left(1-\exp(-g^{\prime}_{i}+\log p_{i})\right)\nabla_{\theta}\log p_{i}$
$=\sum_{i=1}^{|\mathcal{T}|}\left(1-p_{i}\cdot\exp(-g^{\prime}_{i})\right)\nabla_{\theta}\log p_{i}.$
This means that instead of updating the policy only for the selected token, in SofT-GRPO, it inherently highlights: (1). Higher importance component ($g^{\prime}_{i}$) in the soft tokens, (2). lower prob ($p_{i}$). For example, for positive instances, SofT-GRPO inherently improves the probability of the main component in each soft-tokens and the low-probability components. For Negative instances, SofT-GRPO inherently gives penalties to the probability of the main component in each soft-tokens and the low-probability components.
Figure 10 verifies this conclusion: the entropy of token probability distributions will converge to nearly 0 with discrete-token GRPO, but keep stable and even improve with SofT-GRPO.
Figure 10: Entropy curve on DeepSeek-R1-Distill-Qwen-7B during training.
Appendix G Baselines & Datasets & Licenses
In our experiments, we evaluate and compare the following model baselines and datasets. For each, we detail the official website and usage license.
G.1 Baselines
We mainly include No-Finetune base LLMs, Discrete-Token GRPO, and the method in Butt et al. (2025) (noted Soft Token) as baselines.
Base LLMs
This paper includes DeepSeek-R1-Distill-Qwen-1.5B, LLaMA-3.2-3B-Instruct, and DeepSeek-R1-Distill-Qwen-7B as base LLMs.
Discrete-Token GRPO
We utilize the latest verl (Sheng et al., 2024) https://github.com/volcengine/verl/tree/main and vLLM rollout to implement discrete-token GRPO with default parameters.
Butt et al. (2025) Soft Token
Due to the requirement of passing d-dimensional inputs $\hat{\boldsymbol{s}}$ between the rollout workers and the verl policy optimization workers, implementing this algorithm requires a high amount of communication between the rollout workers and the policy update workers. So, we only report the results in (Butt et al., 2025) for comparison in Table 4.
G.2 Datasets
This paper covers five in-domain numerical reasoning datasets (i.e., AIME2024, AIME2025, AMC23, MATH-500, and GSM8K (Cobbe et al., 2021)), one out-of-domain scientific reasoning dataset GPQA-Diamond (Rein et al., 2024), and two out-of-domain code reasoning datasets (i.e., HumanEval and MBPP). These datasets are provided in https://github.com/eric-ai-lab/Soft-Thinking.
G.3 Inference Framework
The inference framework of our implementation is built on SGLang (Zheng et al., 2024), maximizing efficiency via continuous batching, RadixAttention for KV cache reuse, and compressed finite state machines for faster structured output decoding. In the standard flow, the scheduler flattens discrete token inputs into continuous tensors for GPU execution; generated tokens are returned to update the Radix tree state for subsequent steps.
SGLang’s optional overlap scheduler pipelines execution by issuing the next batch ahead via negative integer slot addresses. Compared to Zhang et al. (2025b), we enabled this for Soft-Thinking by storing generated soft probabilities and indices in VRAM-allocated future maps. A custom kernel resolves the slot addresses to fetch these states just-in-time. This design should be faster by eliminating PCIe bottlenecks, allowing non-blocking scheduling with data kept entirely on the GPU.
G.4 Licenses
For Base LLM, Dataset, and frameworks, we list their Licenses in Table 10.
Table 10: A summary of licenses.
Resources Type License URL
DeepSeek-R1-Distill-Qwen-1.5B Base LLM MIT License https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Qwen-1.5B
LLaMA-3.2-3B-Instruct Base LLM Llama 3.2 License https://huggingface.co/meta-llama/Llama-3.2-3B-Instruct
DeepSeek-R1-Distill-Qwen-7B Base LLM MIT License https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Qwen-7B
verl RL-framework Apache-2.0 license https://github.com/volcengine/verl
verl-0.4.x RL-framework Apache-2.0 license https://github.com/volcengine/verl/tree/v0.4.x
SGLang for soft-thinking Inference-framework MIT License https://github.com/eric-ai-lab/Soft-Thinking
AIME2024, AIME2025, GSM8K Dataset MIT License https://github.com/eric-ai-lab/Soft-Thinking
AMC23, MATH-500 Dataset Available Online https://github.com/eric-ai-lab/Soft-Thinking
HumanEval, GPQA-Diamond Dataset MIT License https://github.com/eric-ai-lab/Soft-Thinking
MBPP Dataset Apache-2.0 license https://github.com/eric-ai-lab/Soft-Thinking
Appendix H Visualization
In this section, in Figure 11 and Figure 12, we provide visualization examples for the outcome of SofT-GRPO on two examples (we show $y_{i}$ numbers that are larger than 0.01 in soft-tokens). Our observations are as follows:
  • Compared to Figure 4 in Zhang et al. (2025b), SofT-GRPO preserves the multiply-path thinking pattern, thus providing better soft-thinking reasoning performances.
  • As discussed in Wu et al. (2025a), SofT-GRPO outcomes can observe possible latent search trees. See the 288th soft-token in Figure 11, it blends two discrete tokens with contrary meanings in its top-2 (No and Yeah, probably with totally different embeddings), representing a more abstract concept than any language token. Moreover, to check whether this pattern has already existed in No-Finetune LLM, we searched their output on GSM8K (a total of 1319 instances), where we could not find this probability distribution. So, we can recognize that this probability distribution is highly possible to emerge after SofT-GRPO.
  • Similarly to Figure 4 in Zhang et al. (2025b), SofT-GRPO can preserve interpretability; the contents in the red box (Okay, so I need to…) can be a good explanation for the path of soft-thinking reasoning.
Figure 11: Visualization: DeepSeek-R1-Distill-Qwen-1.5B + SofT-GRPO on the 4-th instance of GSM8K.
Figure 12: Visualization: DeepSeek-R1-Distill-Qwen-1.5B + SofT-GRPO on the 4-th instance of MATH-500.